### v1

# Low-Rank Quantization-Aware Training for LLMs

Abstract
Large language models (LLMs) are omnipresent, however their practical deployment is challenging due to their ever increasing computational and memory demands. Quantization is one of the most effective ways to make them more compute and memory efficient. Quantization-aware training (QAT) methods, generally produce the best quantized performance, however it comes at the cost of potentially long training time and excessive memory usage, making it impractical when applying for LLMs. Inspired by parameter-efficient fine-tuning (PEFT) and low-rank adaptation (LoRA) literature, we propose LR-QAT – a lightweight and memory-efficient QAT algorithm for LLMs. LR-QAT employs several components to save memory without sacrificing predictive performance: (a) low-rank auxiliary weights that are aware of the quantization grid; (b) a downcasting operator using fixed-point or double-packed integers and (c) checkpointing. Unlike most related work, our method (i) is inference-efficient, leading to no additional overhead compared to traditional PTQ; (ii) can be seen as a general extended pretraining framework, meaning that the resulting model can still be utilized for any downstream task afterwards; (iii) can be applied across a wide range of quantization settings, such as different choices quantization granularity, activation quantization, and seamlessly combined with many PTQ techniques. We apply LR-QAT to the LLaMA-2/3 and Mistral model families and validate its effectiveness on several downstream tasks. Our method outperforms common post-training quantization (PTQ) approaches and reaches the same model performance as full-model QAT at the fraction of its memory usage. Specifically, we can train a 7B LLM on a single consumer grade GPU with 24GB of memory.
1 Introduction
In recent years, large language models (LLMs) have emerged as a powerful tool for a plethora of natural language processing tasks. As these models continue to grow in size and capability, addressing their ever increasing computational and memory demands becomes crucial for practical deployment, especially when considering resource-constrained edge devices.
One of the most effective methods to tackle this problem is neural network quantization, which uses low-bit precision for weight and activation tensors. While recent post-training quantization (PTQ) methods can help with decreasing the model size and improving the computational efficiency of LLMs, they typically lead to subpar performance, especially in the case of low-bit ($\leq 4$) quantization. Quantization-aware training (QAT), conversely, yields significantly better model performance compared to PTQ. However, due to extreme model sizes of modern LLMs, using traditional QAT is very computationally expensive and requires a prohibitively high GPU memory usage, making it impractical.
Inspired by parameter-efficient fine-tuning (PEFT) and low-rank adaptation (LoRA) literature, we propose Low-Rank Quantization-Aware Training (LR-QAT) – a lightweight memory-efficient and inference-efficient QAT algorithm for LLMs. LR-QAT reduces the memory requirements of training a 7B LLM from >70GB of GPU memory to <21GB without degrading the predictive performance compared to traditional full-model QAT, making it possible to train such models on a single consumer grade GPU. Unlike most related work that combines low-rank adaptation with quantization, our method is also inference-efficient. After the training is complete, the auxiliary matrices are naturally absorbed into the quantized weight tensor without loss of accuracy and no extra overhead at inference time. LR-QAT is positioned as a general extended pretraining method, as opposed to being strictly a fine-tuning method – the resulting model is a low-bit general pretrained LLM, that can still be utilized for any task afterwards. If needed, our resulting low-bit pretrained LLM can be fine-tuned on specific downstream tasks or used with multiple LoRA adapters for rapid switching between various tasks.
LR-QAT introduces and combines several innovations designed to reduce memory use without sacrificing model performance: (1) A form of QAT with low-rank reparameterization, in which we place the low-rank weights in the integer domain to ensure they align with the quantization grid of the pretrained weights. This allows for seamless fusion during inference into a single low-bit integer matrix. (2) A downcasting operator that represents the frozen pretrained weights as low-bit INT-$b$ ($b\leq 4$) double-packed into INT8 or as fixed-point values stored in INT8. (3) Finally, we combine the proposed quantization formulation with gradient checkpointing to avoid aggressive memory spikes from storing some of the intermediate results in memory for the backward pass.
We apply LR-QAT to the LLaMA-2/3 and Mistral model families and demonstrate its effectiveness on several general language modeling datasets and zero-shot evaluation on some of the common reasoning downstream tasks. Our method outperforms recent post-training quantization (PTQ) approaches and reaches the same predictive performance as full-model QAT at the fraction of its memory usage. Finally, our method can be applied across a wide range of quantization settings (per-channel or per-block weight quantization, activation quantization, etc.) and can be combined with most of other PTQ techniques.
2 Background and related work
Figure 1: Left: Visual illustration of our proposed LR-QAT for memory efficient QAT. Right: Memory requirements for various QAT techniques on LLaMA-2 7B (assuming batch size 1 and a sequence length of 1024).
Neural network quantization is one of the most powerful ways to reduce model footprint, data transfer and compute requirements. By quantizing a model, high bit-width floating point weights and activations can be represented using low-bit numbers. On top of that, by using low-bit fixed-point representations, such as INT8, one can further reduce energy consumption since the fixed-point operations are more efficient than their floating-point counterparts. Quantizing to 8 bits or lower, however, typically introduces quantization noise in the model, resulting in a potential drop in accuracy/perplexity.
In this section, we provide a brief overview of uniform affine quantization and a summary of recent methods for LLM quantization. We will discuss some of the trade-offs of those techniques. Finally, we touch upon the challenges of LLM quantization and some of the limitations of current approaches.
Uniform affine quantization
We use the following definition of the quantization function:
$\widehat{\mathbf{x}}:=q\left(\mathbf{x};\,s,z,b\right)=s\cdot\vphantom{\Bigg{(}}\Big{(}\,\smash{\underbrace{\operatorname{clip}\!\left(\left\lfloor{\frac{\mathbf{x}}{s}}\right\rceil+z;-2^{b-1},2^{b-1}-1\right)}_{\text{\normalsize$=:\mathbf{x}_{\mathbb{Z}}$}}}-z\Big{)},$1)
[…]
Post-training quantization methods
Post-training quantization (PTQ) algorithms take a pretrained high precision (FP32 / FP16 / BF16) network and convert it directly into a fixed-point network without the need for the original training pipeline [2, 8, 11, 26, 33, 38, 45, 47, 48, 68]. These methods are either data-free or only require a small calibration dataset and are generally quite easy to use. Having almost no hyperparameter tuning makes them usable via a single API call as a black-box method to quantize a pretrained neural network in a computationally efficient manner.
Post-training quantization of LLMs is a challenging task due to presence of numerical outliers in weights and activations [6, 7, 32, 14, 55]. Existing LLM PTQ methods can be broadly categorized into weights-only quantization and weight-activation quantization.
Weights-only quantization focuses on converting weights to low-bit values. For instance, GPTQ [18] employs second-order information to iteratively round grouped weights and correct the quantization error in the remaining groups. SpQR [15], AWQ [39] and OWQ [34] emphasize the importance of so-called “salient” weights that correspond to high-magnitude activations. Other recent W-only methods include [29, 36, 44, 9].
Weight-activation quantization compresses both weights and activations. SmoothQuant [62], LLM.int8() [14] and Outlier Suppression [59] achieve W8A8 quantization by managing activation outliers. LLM.int8() uses mixed-precision decomposition, while the other two employ channel-wise scaling. OmniQuant [52] modulates the extreme values of weights by optimizing the clipping threshold and shifts the challenge of quantization from activations to weights by employing the learnable equivalent transformation. Some of the other recent W&A PTQ methods are [35, 41, 60, 65, 56, 64, 40].
Quantization-aware training methods
Quantization-aware training (QAT) methods [4, 17, 21, 27, 33] simulate quantization during training, allowing the model to find more optimal solutions compared to PTQ approaches. However, better accuracy/perplexity comes at the cost of neural network training, i.e., longer training times, increased memory usage, need for labeled data and hyperparameter search.
The excessive training cost and memory usage of traditional QAT methods make them unsuitable for quantizing modern LLMs. A few works that apply QAT to LLMs include LLM-QAT [42] that combine QAT with data-free knowledge distillation, and EdgeQAT [53] that only considers tiny (sub 100M parameter) language models.
Low-rank adapters for fine-tuning
Low-rank adaptation (LoRA) [24] is a parameter efficient fine-tuning (PEFT) method that reduces memory requirements compared to standard training. LoRA freezes the pretrained weights $\bm{W}=\bm{W_{0}}$, and only trains a small set of low-rank trainable parameters, often termed adapters. Given a linear projection $\mathbf{y}=\bm{W}\mathbf{x}$ with $\bm{W}\in\mathbb{R}^{m\times k}$, LoRA computes
$\mathbf{y}=\bm{W}\mathbf{x}+\frac{\alpha}{r}\bm{A}\bm{B}\mathbf{x},$ (2)
where $\bm{A}\in\mathbb{R}^{m\times r}$, $\bm{B}\in\mathbb{R}^{r\times k}$, $r<\min\left\{m,k\right\}$ – rank, and $\alpha$ is a scalar that is constant in $r$. The benefits of LoRA are that it is much cheaper and often performs on par with or better than full fine-tuning. Additionally, the fine-tuned (floating-point) model can be deployed without extra cost, as the low-rank matrices can be fused into the pretrained weights after fine-tuning ($\bm{W}:=\bm{W_{0}}+\frac{\alpha}{r}\bm{A}\bm{B}$).
Naturally, there have been several works that explored the combination of LoRA and quantization. QLoRA [16] quantizes the pretrained weights to 4 bit using NF4 format and dequantizes them in the forward pass to further reduce fine-tuning memory footprint. QA-LoRA [63] uses INT4 quantization and introduces group-wise operators to enable quantization during inference stage. LoftQ [37] proposed an iterative SVD-based procedure for initializing $\bm{A}$, $\bm{B}$ that yields faster fine-tuning convergence when used together with low-bit quantization. LQ-LoRA [20] further extends initialization technique from LoftQ to mixed precision and data aware cases. Other recent works include [28, 67].
Finally, the closest work to ours is PEQA [31], that attempts to combine the benefits of inference-efficiency of QAT together with memory-efficiency of PEFT methods. However, their approach is different since they focus on a task-specific fine-tuning as opposed to being a general extended pretraining method. In addition to that, PEQA has significantly less degrees of freedom compared to our method, leading to a subpar performance.
Motivation
PTQ, while generally fast and simple, suffers from limited performance in low-bit scenarios. Although QAT methods still perform well in low-bit regimes, their high training costs and memory usage make them impractical for LLMs.
LoRA-based methods address memory issues for efficient fine-tuning. However, in most cases they don’t consider efficient inference. The adapters $\bm{A}$, $\bm{B}$ are typically stored in higher precision format such as BF16 and at inference they dequantize the low-bit integer matrix $\bm{W}_{\mathbb{Z}}$ to the same data format, resulting in runtime overhead.
Simply quantizing adapters after training will lead to a different quantization grid compared to $\bm{W}$, and quantizing them using the same quantization grid as $\bm{W}$ will lead to high error. QA-LoRA is the only work we are aware of that attempts to fuse auxiliary LoRA weights back into the frozen $\bm{W}_{\mathbb{Z}}$. Yet, their method is designed to only work with group-wise quantization with high number of groups (a small group size of $32$). In addition to that, QA-LoRA and most of LoRA-based methods combine their proposed technique with the task-specific fine-tuning, whereas we propose LR-QAT as an extended pretraining method.
We are inspired by LoRA-based methods to make QAT more memory- and runtime-efficient. In addition to that, our goal is to design a method that is inference-efficient, i.e. where the low-rank adapters can be fused back into a low-bit integer matrix $\bm{W}_{\mathbb{Z}}$ without any loss of accuracy/perplexity, yielding PTQ level of inference efficiency. Contrary to QA-LoRA [63], we are not relaxing the quantization constraints – our method is applicable at any weight quantization granularity. Finally, we see our method as a general extended pretraining framework. The resulting model can afterwards still be used on any task. We summarize different trade-offs for the discussed techniques in Table 1.
Method Accuracy Memory efficiency Inference efficiency
PTQ
(Full-model) QAT
LoRA / PEFT
[…]
Table 1: Comparison between existing approaches and the proposed method.
3 Method
We now discuss the components of LR-QAT followed by a formal definition of LR-QAT.
QAT with low-rank adapters
Let’s recall how traditional QAT works. Given a linear layer with a weight matrix $\bm{W}\in\mathbb{R}^{m\times k}$ and assuming $b$-bit symmetric uniform affine quantization, the quantization is simulated as follows:
$\widehat{\bm{W}}:=\mathbf{s}\cdot\operatorname{clip}\left(\left\lfloor{\frac{\bm{W}}{\mathbf{s}}}\right\rceil,-2^{b-1},2^{b-1}-1\right),$ (3)
where weights $\bm{W}$ (often termed “shadow weights”) are trainable parameters and the quantization scale $\mathbf{s}$ can be either fixed or also learned. To be able to backpropagate through round-to-nearest operation in (3), it is common to use straight-through estimator (STE, Bengio et al. 3), where it is assumed that $\frac{\partial\!\left\lfloor{t}\right\rceil}{\partial t}=1$. When applied to LLMs, it is straightforward to see that this procedure is very expensive: we have to learn a comparable number of parameters that was used for pretraining, leading to excessive memory usage.
To make this approach more practical we freeze the pretrained weights $\bm{W}$ (denote $\bm{W_{0}}$) and introduce low-rank adapters $\bm{A}\in\mathbb{R}^{m\times r}$, $\bm{B}\in\mathbb{R}^{r\times k}$, $r\ll\min\left\{m,k\right\}$. We have to be careful where exactly those adapters are placed. As discussed in Section 2, after the training is complete, we want $\bm{A}$ and $\bm{B}$ to be seamlessly integrated into a single $b$-bit integer matrix $\bm{W}_{\mathbb{Z}}$ without loss of accuracy to facilitate the efficient inference. To accommodate that, we put the auxiliary matrices inside the rounding operator as follows
$\widehat{\bm{W}}:=\mathbf{s}\cdot\operatorname{clip}\left(\left\lfloor{\frac{\bm{W_{0}}}{\mathbf{s}}+\frac{\alpha}{r}\bm{A}\bm{B}}\right\rceil,-2^{b-1},2^{b-1}-1\right),$ (4)
where we are using STE assumption for the rounding operation to compute the gradients of the loss w.r.t. $\bm{A}$, $\bm{B}$ and $\mathbf{s}$. We further employ a scaling factor $\alpha/r$ used in LoRA [24] to reduce the need to retune hyperparameters as we vary the rank $r$. After training is complete, (4) can be represented as regular fixed point tensor, $\widehat{\bm{W}}=\mathbf{s}\cdot\bm{W}_{\mathbb{Z}}$, without any extra effort or loss of accuracy and therefore enabling efficient inference without any extra overhead. Note that this is different to most of the literature, such as QLoRA [16], where adapters are placed outside of the quantization function (such as $\mathbf{y}=\widehat{\bm{W}}\mathbf{x}+\bm{A}\bm{B}\mathbf{x}$) and typically stored in higher precision formats such as BF16.
Downcasting operator
The formulation (4) is already significantly more memory efficient compared to standard full-model QAT (3). We don’t need to compute neither gradients w.r.t. weights $\bm{W}$ nor the respective first or second-order momentum terms for Adam-based optimizers, and only need to do so for the auxiliary matrices $\bm{A}$ and $\bm{B}$, which is noticeably more affordable provided $r\ll\min\left\{m,k\right\}$.
Given that the weight matrix $\bm{W_{0}}$ is frozen, the next natural step to further reduce the memory utilization is to store it in a lower-precision format. One could apply downcasting of $\bm{W_{0}}$ directly in (4), however notice that those weights are divided by the scale $\mathbf{s}$ at every forward pass, which generally has to be stored in a high-precision format to guarantee stable training. To simplify further, we propose the following variant of low-rank QAT:
$\widehat{\bm{W}}:=\mathbf{s}\cdot\operatorname{clip}\left(\left\lfloor{\frac{\bm{W_{0}}}{\mathbf{s_{0}}}+\frac{\alpha}{r}\bm{A}\bm{B}}\right\rceil,-2^{b-1},2^{b-1}-1\right),$ (5)
where we use the initial scale [^1] A frozen scale obtained after initial range estimation before the training begins. $\mathbf{s_{0}}$ instead of learned scale $\mathbf{s}$ inside the rounding operator, and the rest is the same as in (4). Now the entire fraction $\bm{W_{0}}/\mathbf{s_{0}}$ is fixed and we can store it in a lower-precision format. Note that the scale $\mathbf{s}$ outside of the clipping operator can still be learned. Empirically, we found that (5) performs consistently on par with or even slightly better compared to (4).
During training the pretrained weights are represented and stored as follows
$\bm{\Phi}_{0}:=\varphi\left(\frac{\bm{W_{0}}}{\mathbf{s_{0}}}\right),$ (6)
where $\varphi\left(\cdot\right)$ – a downcasting operator that encapsulates a choice of different numeric formats or other preprocessing computations. In the simplest form, $\varphi\left(\cdot\right)$ would cast the input to one of pre-existing floating-point formats, such as FP16, BF16, FP8 etc.
Inspired by traditional fixed point quantization, we also explore integer representations for $\varphi\left(\cdot\right)$. Specifically, $\varphi\left(x\right)=\operatorname{clip}\left(\left\lfloor{x}\right\rceil,-2^{b-1},2^{b-1}-1\right)$ corresponds to a standard $b$-bit integer quantization and can be stored as INT-$b$ number. We denote this approach $\varphi={\text{{\small{INT-}}$b$}}$ for brevity. In addition to that, in case of low-bit quantization ($b\leq 4$), which is a primary focus of this work, two INT-$b$ numbers can be double-packed into a single INT8 number, leading to further memory savings. This is helpful because most of the common deep learning frameworks like PyTorch, at the time of writing this paper, don’t natively support low-bit formats such as INT4 yet.
[…]
We address this problem in two distinct ways: We adapt and experiment with a variant of SVD-based initialization for low-rank matrices $\bm{A}$, $\bm{B}$ proposed in [37] before we apply a downcasting operator to $\bm{W_{0}}/\mathbf{s_{0}}$, to capture some of the information of it’s fractional part. With this approach we can still employ a double-packing since we are still using $\varphi={\text{{\small{INT-}}$b$}}$.
Another way is to use INT8 storage type, allocate $b$ bits to represent the integer part as before, but utilize the remaining $8-b$ bits for storing the approximate fractional part ($2\leq b\leq 7$). In other words, we represent $\bm{\Phi}_{0}$ using fixed-point numbers. Specifically, assuming the the rest of the computation is performed in BF16, we define the downcasting and the corresponding upcasting operators as follows:
[…]
A fixed-point number where $n$ bits are used for the integer part of the value and $m$ bits are used for the fractional part are commonly denoted [50] as Qn.m. For brevity, we will refer to (7) as $\varphi={\text{{\small{Q$b$.$(8-b)$}}}}$. Specifically, in this work we will be mainly focusing on $b\in\left\{3,4\right\}$, which corresponds to Q3.5 and Q4.4, respectively.
Gradient checkpointing
Note that both in the original LoRA paper [24] and in the related work like QLoRA [16], there is no need to compute the product $\bm{A}\bm{B}$. Instead, those matrices are multiplied with the activations $\mathbf{x}$ as $\bm{A}\left(\bm{B}\mathbf{x}\right)$. However, we do compute a product $\bm{A}\bm{B}$ in (5), and in the naive implementation of our method, this product together with the results of some intermediate computations (e.g., after rounding and clipping) will be automatically kept in memory for the backward pass, leading to increased memory usage. To prevent this, we employ gradient checkpointing [10] on (5). In other words, we recompute the quantizer function in the backward pass, leading to a slight runtime overhead but avoiding significantly increased memory usage.
LR-QAT
Using the components described above, we define LR-QAT for a single layer with a (pretrained) weight matrix $\bm{W_{0}}$ as follows
$\widehat{\bm{W}}:=\mathbf{s}\cdot\operatorname{clip}\left(\left\lfloor{\bm{\Phi}_{0}+\frac{\alpha}{r}\bm{A}\bm{B}}\right\rceil,-2^{b-1},2^{b-1}-1\right),$ (8)
where $\mathbf{s}$ – trainable or frozen quantization scale with the initial value of $\mathbf{s_{0}}$, $\bm{A}$, $\bm{B}$ – trainable rank $r$ auxiliary matrices, $\bm{\Phi}_{0}:=\varphi\left(\bm{W_{0}}/\mathbf{s_{0}}\right)$ – frozen representation of the original pretrained weights, and $\varphi$ is the downcasting operator. To avoid excessive memory allocation for the results of intermediate computations in (8) involving the product $\bm{A}\bm{B}$, we apply checkpointing on $\widehat{\bm{W}}$. After the training is complete, low-rank adapters are naturally integrated into a single integer matrix $\bm{W}_{\mathbb{Z}}=\operatorname{clip}\left(\cdots\right)$ without loss of accuracy. Note, while we presented our method for symmetric quantization which is commonly used for weights [49], it can equally be applied for asymmetric quantization by adding a zero offset $z$ outside the rounding operation as shown in (1).
4 Experiments
We assess the effectiveness of LR-QAT by conducting experiments on LLaMA-2 7B/13B [58], LLaMA-3 8B [1] and Mistral-0.1 7B [30]. To compare to relevant literature with weight-activation quantization, we further experiment with LLaMa-1 7B [57]. We first explore the impact of the choice of rank $r$, a downcasting operator $\varphi\left(\cdot\right)$, and the initialization of auxiliary matrices $\bm{A}$, $\bm{B}$. We then compare our method in terms of accuracy to standard full-model QAT, other baselines, and the related work. All detailed hyperparameters of our experiments are in Appendix A.
Quantization
We experiment with both weight-only and weight-activation quantization. The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’). We quantize all linear layers, except the classification head. In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14]. Following OmniQuant [52] we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [42].
Evaluation
Following the previous work [18, 62, 52, 42], we evaluate quantized models by reporting the perplexity of language generation on WikiText-2 [46], using a sequence length of $2048$. We also report zero-shot accuracy on a set of common sense reasoning tasks including BoolQ [12], PIQA [5], Winogrande [51], ARC [13], and HellaSwag [66]. For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
Datasets and training
We train on a small subset of SlimPajama [54], which is a close open-source replica of the dataset used for pretraining LLaMA models. For weight-only quantization results, we train for $10^{4}$ steps using batch size $32$ and a maximum sequence length of $1024$. For ablation studies and weight-activation quantization results, we use shorter training of $10^{3}$ steps. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set (512 sequences).
Baselines
We compare with round-to-nearest quantization (RTN), where we set the ranges based on minimizing the $L^{p}$-norms between quantized and unquantized weights and report the best performing configuration. We also use that as initialization for LR-QAT. For weight-only quantization, we compare with GPTQ [18], AWQ [39], OmniQuant [52], and our implementation of PEQA [31], where we use symmetric weight quantization and following the same experimental setup and best RTN initialization, for a fair comparison. For weight-activation quantization, we compare to RTN, SmoothQuant [62] and LLM-QAT [42]. Following [42], we compare to them in several different settings, where the weights, activations and KV cache values are quantized to different levels (denoted as W-A-KV).
4.1 The impact of rank $r$
{floatrow}
[…]
$\varphi\left(\cdot\right)$ dtype $\bm{A}$, $\bm{B}$ init. W4 @ pc W3 @ pc
WT-2 $\downarrow$ zero-shot $\uparrow$ WT-2 $\downarrow$ zero-shot $\uparrow$
FP32 FP32 LoRA 5.69 69.28 6.21 66.62
FP16 FP32 LoRA $+$0.00 --0.13 $+$0.01 --0.01
BF16 FP32 LoRA --0.01 $+$0.11 $+$0.01 $+$0.45
Q4.4 / Q3.5 FP32 LoRA --0.01 $+$0.16 $+$0.01 $+$0.31
Q4.4 / Q3.5 BF16 LoRA --0.01 $+$0.15 $+$0.01 $+$0.31
INT-4 / INT-3 FP32 LoRA $+$0.02 --0.04 $+$1.5e3 --30.30
INT-4 / INT-3 FP32 LoftQ ($T=1$) $+$0.14 --1.02 $+$0.18 $+$0.26
INT-4 / INT-3 FP32 LoftQ ($T=64$) $+$0.40 --1.40 $+$1.37 --2.01
[…]
[figure]
[…]
We investigate the effect of different values of rank $r$ of the auxiliary matrices $\bm{A}$ and $\bm{B}$ and present results in Figure 2. Increasing the rank from $1$ to $32$ leads to progressively slightly better performance, excluding one outlier. The fact that using $r>32$ doesn’t lead to further improvement in perplexity is likely because of the limited number of training steps we used for this experiment ($10^{3}$), and that more steps might be needed for the procedure to fully converge. Interestingly, a rank $r$ as small as 1 already performs really well. We hypothesize that this is the case because of the following. Even though $\operatorname{rank}\left(\bm{A}\bm{B}\right)=1$, by applying a low-rank approximation inside the rounding and clipping operators in (8), this can overall lead to a high-rank perturbation to the original weights $\bm{\Phi}_{0}$ (in the integer domain). Going forward, we use $r=32$ in all our experiments [^2] This amounts to only 1.2% of the total number of parameters for 7B LLaMA model..
4.2 The impact of downcasting operator $\varphi\left(\cdot\right)$ and initialization
We study the effect of several choices of the downcasting operators discussed in Section 3 and summarize results in Table 2. We can see that by going from FP32 to BF16, and finally to an 8-bit fixed-point representation of $\bm{\Phi}_{0}$, aside from memory savings we also maintain the same WikiText-2 perplexity and even slightly improve zero-shot accuracy. The latter is likely due to a slight regularization effect caused by the fact that we discard some of the information in the fractional part in $\bm{W_{0}}/\mathbf{s_{0}}$, some of which might be noise. One step further, however, while $\varphi{}={\text{{\small{INT-$b$}}}}$ still leads to a good model performance in the case of 4-bit weight quantization, it completely breaks for W3.
[…]
Finally, when using the fixed point representation for $\bm{\Phi}_{0}$, we still maintain the same model performance by switching the compute data type [^3] A data type used for activations, gradients, and frozen parameters. from FP32 to BF16, where the latter is what is commonly used for LLMs.
4.3 Comparison with full-model QAT
Table 3: Comparison of the proposed method with the vanilla (full-model) QAT on LLaMA-2 7B with W4 per-channel quantization.
Method GPU memory, GB WikiText-2 perplexity $\downarrow$ Zero-shot avg. accuracy $\uparrow$
Full model QAT 71.9 5.74 69.29
LR-QAT 20.5 5.68 69.43
[…]
4.4 Main results
Table 4: Weight-only quantization results for LLaMA-2/3 and Mistral models. We report WikiText-2 test set perplexity (lower is better) and zero-shot average accuracy (higher is better). Models marked ‘L2’, ‘L3’, and ‘M’ denote LLaMA-2, LLaMA-3, and Mistral, respectively. Numbers marked in bold are SOTA or on par (within $0.05$). §Uses asymmetric weight quantization.
Bitwitdth Method WikiText-2 Perplexity $\downarrow$ Zero-shot avg. accuracy $\uparrow$
L2-7B L2-13B L3-8B M-7B L2-7B L2-13B L3-8B M-7B
FP16 5.47 4.88 6.14 5.25 70.47 73.18 74.22 75.69
W4 @ pc RTN 6.14 5.21 7.53 5.91 68.88 71.73 72.19 73.44
GPTQ§ 5.83 5.13 - - - - - -
AWQ 6.15 5.12 - - - - - -
OmniQuant§ 5.74 5.02 - - - - - -
PEQA (our impl.) 5.71 5.03 7.51 5.56 69.23 72.51 72.79 73.73
LR-QAT (ours) 5.66 5.04 6.78 5.46 69.95 73.13 73.35 73.67
W3 @ pc RTN 26.73 8.71 34.10 9.49 43.87 55.01 47.46 64.58
GPTQ§ 8.37 6.44 - - - - - -
AWQ 24.00 10.45 - - - - - -
OmniQuant§ 6.58 5.58 - - - - - -
PEQA (our impl.) 6.45 5.73 27.45 6.51 65.44 69.81 50.28 71.02
LR-QAT (ours) 6.13 5.54 8.12 6.03 67.66 71.22 70.46 71.87
W4 @ g128 RTN 5.78 5.04 6.96 5.49 69.75 72.94 72.04 75.07
GPTQ§ 5.61 4.98 - - - - - -
AWQ 5.62 4.97 - - - - - -
OmniQuant§ 5.58 4.95 - - - - - -
PEQA (our impl.) 5.67 5.02 6.89 5.48 69.64 72.80 72.99 73.34
LR-QAT (ours) 5.59 4.97 6.62 5.36 69.88 72.89 73.72 75.18
W3 @ g128 RTN 7.61 6.14 15.11 6.77 63.20 66.81 57.74 69.35
GPTQ§ 6.29 5.42 - - - - - -
AWQ 6.24 5.32 - - - - - -
OmniQuant§ 6.03 5.28 - - - - - -
PEQA (our impl.) 6.05 5.58 9.64 5.85 68.10 70.29 67.19 72.21
LR-QAT (ours) 5.99 5.53 7.74 5.81 67.98 70.54 70.48 71.44
Weight-only quantization
We summarize our results in Table 4. As we can see, in almost most cases LR-QAT outperforms or is on par with prior weight-only quantization methods across various LLM families and quantization settings, including both per-channel and group-wise quantization. In a few cases, especially in case of group-wise quantization our method did not outpeform OmniQuant. However, OmniQuant uses asymmetric quantization which provides extra degrees of freedom compared to symmetric quantization, which are very helpful in the case of low-bit quantization. In practice, however, symmetric weight quantization yields more efficient inference [49]. Additionally, techniques like OmniQuant and related techniques are orthogonal to our method and can be used as as initialization of LR-QAT.
Weight-activation quantization
Table 5: Weight and activation quantization results for LLaMA-1 7B. We report WikiText-2 test set perplexity (lower is better) and zero-shot accuracy of 6 tasks (higher is better). Numbers marked in bold are SOTA. *Evaluated with a maximum sequence length of 1024.
Bitwitdth (W-A-KV) Method WikiText-2 $\downarrow$ Zero-shot avg. accuracy $\uparrow$
BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
FP16 5.68 75.05 79.16 70.01 72.85 44.80 76.21 69.68
4-8-8 RTN 3.4e4 45.87 49.13 48.46 25.51 28.50 26.01 37.25
SmoothQuant - 71.00 76.00 66.00 67.40 42.80 67.80 65.17
LLM-QAT 11.2* 74.60 77.50 67.70 70.20 45.60 73.50 68.18
LR-QAT (ours) 5.97 72.97 77.86 68.98 69.91 41.30 74.35 67.56
4-8-4 RTN 3.5e4 40.37 49.46 49.88 26.18 28.50 26.11 36.75
SmoothQuant - 54.70 55.40 51.50 43.90 27.70 38.90 45.35
LLM-QAT 11.6* 69.50 75.40 64.60 66.00 43.80 69.20 64.75
LR-QAT (ours) 6.01 75.60 78.62 69.93 69.74 41.04 74.51 68.24
We present our results for weight-activation quantization applied to LLaMA-1 7B in Table 5. LR-QAT consistently outperforms the PTQ baselines and is on par or better compared to LLM-QAT. This demonstrates that the proposed method is readily applicable not only to weight-only quantization but also weight-activation quantization, a setting that allows for a very efficient inference using fixed-point arithmetic. Finally, our method can still be combined with most of the related PTQ methods including OmniQuant that shift the difficulty of activation quantization to weight quantization, and will likely lead to even better results.
5 Discussion
Limitations
A core assumption of LR-QAT is that a low-rank approximation can compensate the introduced quantization noise. While quantization noise follows a random uniform distribution and is theoretically not low-rank, our results and several prior works [16, 63, 31] suggest that in an end-to-end training setup, low-rank approaches can effectively compensate quantization noise. In our work, we demonstrated the effectiveness of LR-QAT for LLMs up to 13B parameters. It is unclear how it scales to significantly larger LLMs, though we do not see any reason why our findings should not hold beyond 13B models. We evaluate LR-QAT as extended pretraining technique with several thousands of iterations. It is unclear how it would perform in case it is used during pretraining for millions of iterations.
Impact
As LR-QAT improves LLM quantization and makes them therefore more efficient, we expect mainly positive outcomes of our work. Efficiently deploying LLMs will help with reducing their high power consumption at inference time. It further helps to move inference from the cloud to edge devices which can overcome potential privacy concerns. In some cases, quantization might lead to biased predictions, see [23] for further discussion.
6 Conclusions
In this paper we propose LR-QAT, a lightweight and memory-efficient QAT algorithm for LLMs which enables training a 7B LLM on a single consumer grade GPU with 24GB of memory. Inspired by PEFT methods, we introduce a low-rank reparameterization that is aware of the quantization grid. We further reduce the memory requirements by introducing a downcasting operator involving fixed-point or double-packed integers, and applying checkpointing. In almost all cases, our method outperforms common PTQ approaches and reaches the same model performance as full-model QAT at the fraction of its memory usage.
Acknowledgements
We would like to thank Tijmen Blankevoort, Paul Whatmough, Jorn Peters, and Ties van Rozendaal for valuable discussions and support.
Appendix A Experimental details
In this section, we list the details related to hyperparameters and other settings used in our experiments. If not stated otherwise, the standard hyperparameters that we use are the following:
Hyperparameter Value
Optimizer AdamW
LR for $\bm{A}$, $\bm{B}$ $10^{-2}$
LR Schedule for $\bm{A}$, $\bm{B}$ Linear with Warmup
LR for Quantization scales $10^{-5}$
LR Schedule for Quantization scales Linear with Warmup
Warmups Steps $10\%$ of Training Steps
Weight Decay $0.0$
Weight Decay for $\bm{A}$, $\bm{B}$ $0.0$
Adam $\beta_{1}$ $0.90$
Adam $\beta_{2}$ $0.95$
Training Steps $10^{4}$
Gradient Clipping (Maximum $L^{2}$-norm) $1.0$
Table 6: Common hyperparameters used for experiments.
Quantization
We experiment with both weight-only and weight-activation quantization. The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’). We always use symmetric uniform affine quantization. We quantize all linear layers, except the classification head. LayerNorm and Embedding layers are always kept at full precision. In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14]. Following OmniQuant [52] we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [42].
Datasets and training
For all our experiments we use the linear with warmup learning rate schedule from HugginFace’s transformers package [61].
To optimize the learnable parameters, we use AdamW optimizer [43] with weight decay set to zero, $\beta_{1}=0.9$ and $\beta_{2}=0.95$. We use a separate learning rate for quantization scales and for low-rank adapters. Quantization scales learning rate is always set to $10^{-5}$, where the learning rate for low-rank adapters is tuned depending on the experiment. For all the weight-only quantization experiments we set to $10^{-2}$, except W4g128 where we found that a smaller learning rate of $10^{-4}$ or $10^{-5}$ is beneficial. For weight-activation quantization experiments we set the learning rate to $10^{-5}$.
We train on a small subset of SlimPajama [54], which is a close open-source replica of the dataset used for pre-training LLaMA models. In all experiments we set $\alpha$ in (8) to $1$. We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set [^4] Specifically, we use the English subset of Wiki40b, https://huggingface.co/datasets/wiki40b, that contains cleaned-up text of English Wikipedia and training/validation splits. (512 sequences).
For weight-only quantization results we train for $10^{3}$ steps, but for LLaMA-v3 and Mistral we only train for $5000$ steps due to constrained resources. In all cases we used a number of warmup steps equal to the $10\%$ of the training steps, a batch size of $32$, and a maximum sequence length of $1024$. For ablation studies we always use shorter training of $1000$ steps, and for weight-activation quantization results we use training of $1000$ steps.
Baselines
We compare with vanilla round-to-nearest quantization (RTN), where we explore several choices of range setting and report the best one based on Wikipedia validation set perplexity, and also use that as initialization for our method. Specifically, we experimented with min-max range estimator and with $L^{p}$-norm range estimator with the following values for $p$: $\{2.0,2.4,3.0,3.5,4.0,5.0\}$.
Computational Resources
Each experiment for which we reported results, was executed on a single NVidia A100 GPU equipped with 80GB of VRAM. LLaMA-2 7B and 13B experiments needed respectively around 3 and 5 days for 10k training steps experiments. Mistral 7B and LLaMA-3 8B needed around 1.6 days for 5k training steps experiments. For obtaining all the results in the paper, including the ablations, we needed 107 GPU days (A100). Including preliminary experiments that did not make it in the final paper and hyperparameter turning we estimate the total compute costs of this research to approximately 500 GPU days.

### v2

@@
 Unlike most related work, our method (i) is inference-efficient, leading to no additional overhead compared to traditional PTQ; (ii) can be seen as a general extended pretraining framework, meaning that the resulting model can still be utilized for any downstream task afterwards; (iii) can be applied across a wide range of quantization settings, such as different choices quantization granularity, activation quantization, and seamlessly combined with many PTQ techniques.
-We apply LR-QAT to the LLaMA-2/3 and Mistral model families and validate its effectiveness on several downstream tasks.
+We apply LR-QAT to LLaMA-2/3 and Mistral model families and validate its effectiveness on several downstream tasks.
 Our method outperforms common post-training quantization (PTQ) approaches and reaches the same model performance as full-model QAT at the fraction of its memory usage.
@@
 (3) Finally, we combine the proposed quantization formulation with gradient checkpointing to avoid aggressive memory spikes from storing some of the intermediate results in memory for the backward pass.
-We apply LR-QAT to the LLaMA-2/3 and Mistral model families and demonstrate its effectiveness on several general language modeling datasets and zero-shot evaluation on some of the common reasoning downstream tasks.
-Our method outperforms recent post-training quantization (PTQ) approaches and reaches the same predictive performance as full-model QAT at the fraction of its memory usage.
-Finally, our method can be applied across a wide range of quantization settings (per-channel or per-block weight quantization, activation quantization, etc.) and can be combined with most of other PTQ techniques.
+We apply LR-QAT to LLaMA-2/3 and Mistral model families and demonstrate its effectiveness on several general language modeling datasets and zero-shot evaluation on some of the common reasoning downstream tasks.
+Our method outperforms recent PTQ approaches and reaches the same predictive performance as full-model QAT at the fraction of its memory usage.
+Finally, our method can be applied across a wide range of quantization settings, including per-channel or per-block weight quantization, activation quantization, and can be combined with most of other PTQ techniques.
 2 Background and related work
-Figure 1: Left: Visual illustration of our proposed LR-QAT for memory efficient QAT.
-Right: Memory requirements for various QAT techniques on LLaMA-2 7B (assuming batch size 1 and a sequence length of 1024).
+Figure 1: Left: A schematic illustration of our proposed LR-QAT.
+$\mathbf{x}$ and $\mathbf{y}$ denote the input and the output of the linear layer.
+Right: Memory requirements for training with various QAT techniques on LLaMA-2 7B (assuming batch size 1, sequence length 1024, $r=32$, and BF16 compute data type).
 Neural network quantization is one of the most powerful ways to reduce model footprint, data transfer and compute requirements.
@@
 LoRA freezes the pretrained weights $\bm{W}=\bm{W_{0}}$, and only trains a small set of low-rank trainable parameters, often termed adapters.
-Given a linear projection $\mathbf{y}=\bm{W}\mathbf{x}$ with $\bm{W}\in\mathbb{R}^{m\times k}$, LoRA computes
-$\mathbf{y}=\bm{W}\mathbf{x}+\frac{\alpha}{r}\bm{A}\bm{B}\mathbf{x},$ (2)
+Given a linear projection $\mathbf{y}=\bm{W_{0}}\mathbf{x}$ with $\bm{W_{0}}\in\mathbb{R}^{m\times k}$, LoRA computes
+$\mathbf{y}=\bm{W_{0}}\mathbf{x}+\frac{\alpha}{r}\bm{A}\bm{B}\mathbf{x},$ (2)
 where $\bm{A}\in\mathbb{R}^{m\times r}$, $\bm{B}\in\mathbb{R}^{r\times k}$, $r<\min\left\{m,k\right\}$ – rank, and $\alpha$ is a scalar that is constant in $r$.
@@
 Naturally, there have been several works that explored the combination of LoRA and quantization.
-QLoRA [16] quantizes the pretrained weights to 4 bit using NF4 format and dequantizes them in the forward pass to further reduce fine-tuning memory footprint.
+QLoRA [16] quantizes the pretrained weights to 4 bit using (a non-uniform) NF4 format and dequantizes them in the forward pass to further reduce fine-tuning memory footprint.
 QA-LoRA [66] uses INT4 quantization and introduces group-wise operators to enable quantization during inference stage.
@@
 Motivation
-PTQ, while generally fast and simple, suffers from limited performance in low-bit scenarios.
+While generally fast and simple, PTQ suffers from limited performance in low-bit scenarios.
 Although QAT methods still perform well in low-bit regimes, their high training costs and memory usage make them impractical for LLMs.
@@
 However, in most cases they don’t consider efficient inference.
-The adapters $\bm{A}$, $\bm{B}$ are typically stored in higher precision format such as BF16 and at inference they dequantize the low-bit integer matrix $\bm{W}_{\mathbb{Z}}$ to the same data format, resulting in runtime overhead.
-Simply quantizing adapters after training will lead to a different quantization grid compared to $\bm{W}$, and quantizing them using the same quantization grid as $\bm{W}$ will lead to high error.
+The adapters $\bm{A}$ and $\bm{B}$ are typically stored in higher precision formats, such as BF16.
+During inference, the frozen low-bit pretrained weights $\bm{W_{0}}$ are dequantized to match the same data format, resulting in runtime overhead.
+Simply quantizing adapters after training will lead to a different quantization grid compared to $\bm{W_{0}}$, and quantizing them using the same quantization grid as $\bm{W_{0}}$ will lead to a high error.
 QA-LoRA is the only work we are aware of that attempts to fuse auxiliary LoRA weights back into the frozen $\bm{W}_{\mathbb{Z}}$.
 Yet, their method is designed to only work with group-wise quantization with high number of groups (a small group size of $32$).
-In addition to that, QA-LoRA and most of LoRA-based methods combine their proposed technique with the task-specific fine-tuning, whereas we propose LR-QAT as an extended pretraining method.
+In addition to that, QA-LoRA and most of LoRA-based methods combine their proposed techniques with the task-specific fine-tuning, whereas we propose LR-QAT as an extended pretraining method.
 We are inspired by LoRA-based methods to make QAT more memory- and runtime-efficient.
@@
 […]
-Table 1: Comparison between existing approaches and the proposed method.
+Table 1: A comparison between existing approaches and the proposed method.
 3 Method
@@
 $\widehat{\bm{W}}:=\mathbf{s}\cdot\operatorname{clip}\left(\left\lfloor{\frac{\bm{W}}{\mathbf{s}}}\right\rceil,-2^{b-1},2^{b-1}-1\right),$ (3)
-where weights $\bm{W}$ (often termed “shadow weights”) are trainable parameters and the quantization scale $\mathbf{s}$ can be either fixed or also learned.
+where weights $\bm{W}$ are trainable parameters and the quantization scale $\mathbf{s}$ can be either fixed or also learned.
 To be able to backpropagate through round-to-nearest operation in (3), it is common to use straight-through estimator (STE, Bengio et al.
@@
 We have to be careful where exactly those adapters are placed.
-As discussed in Section 2, after the training is complete, we want $\bm{A}$ and $\bm{B}$ to be seamlessly integrated into a single $b$-bit integer matrix $\bm{W}_{\mathbb{Z}}$ without loss of accuracy to facilitate the efficient inference.
+As discussed in Section 2, after the training is complete, we want $\bm{A}$ and $\bm{B}$ to be seamlessly integrated into a single $b$-bit integer matrix $\bm{W}_{\mathbb{Z}}$ without loss of accuracy to facilitate efficient inference.
 To accommodate that, we put the auxiliary matrices inside the rounding operator as follows
@@
 After training is complete, (4) can be represented as regular fixed point tensor, $\widehat{\bm{W}}=\mathbf{s}\cdot\bm{W}_{\mathbb{Z}}$, without any extra effort or loss of accuracy and therefore enabling efficient inference without any extra overhead.
-Note that this is different to most of the literature, such as QLoRA [16], where adapters are placed outside of the quantization function (such as $\mathbf{y}=\widehat{\bm{W}}\mathbf{x}+\bm{A}\bm{B}\mathbf{x}$) and typically stored in higher precision formats such as BF16.
+Note that this is different to most of the literature, such as QLoRA [16], where adapters are placed outside of the quantization function (such as $\mathbf{y}=\widehat{\bm{W}}\mathbf{x}+\bm{A}\bm{B}\mathbf{x}$) and are typically stored in higher precision formats such as BF16.
 Downcasting operator
@@
 Given that the weight matrix $\bm{W_{0}}$ is frozen, the next natural step to further reduce the memory utilization is to store it in a lower-precision format.
-One could apply downcasting of $\bm{W_{0}}$ directly in (4), however notice that those weights are divided by the scale $\mathbf{s}$ at every forward pass, which generally has to be stored in a high-precision format to guarantee stable training.
-To simplify further, we propose the following variant of low-rank QAT:
+One could directly apply downcasting to $\bm{W_{0}}$ in (4).
+However, it’s important to note that these weights are divided by the scale $\mathbf{s}$ during every forward pass.
+To ensure stable training, the scale generally needs to be stored in a high-precision format.
+Therefore, to simplify further, we propose the following variant of low-rank QAT:
 $\widehat{\bm{W}}:=\mathbf{s}\cdot\operatorname{clip}\left(\left\lfloor{\frac{\bm{W_{0}}}{\mathbf{s_{0}}}+\frac{\alpha}{r}\bm{A}\bm{B}}\right\rceil,-2^{b-1},2^{b-1}-1\right),$ (5)
@@
 $\bm{\Phi_{0}}:=\varphi\!\left(\frac{\bm{W_{0}}}{\mathbf{s_{0}}}\right),$ (6)
-where $\varphi\left(\cdot\right)$ – a downcasting operator that encapsulates a choice of different numeric formats or other preprocessing computations.
+where $\varphi\!\left(\cdot\right)$ is a downcasting operator that encapsulates a choice of different numeric formats or other preprocessing computations.
 In the simplest form, $\varphi\!\left(\cdot\right)$ would cast the input to one of pre-existing floating-point formats, such as FP16, BF16, FP8 etc.
@@
 In other words, we represent $\bm{\Phi_{0}}$ using fixed-point numbers.
-Specifically, assuming the the rest of the computation is performed in BF16, we define the downcasting and the corresponding upcasting operators as follows:
+Assuming the rest of the computation is performed in BF16, we define the downcasting and the corresponding upcasting operators as follows:
 […]
@@
 For brevity, we will refer to (7) as $\varphi={\text{{\small{Q$b$.$(8-b)$}}}}$.
-Specifically, in this work we will be mainly focusing on $b\in\left\{3,4\right\}$, which corresponds to Q3.5 and Q4.4, respectively.
+In this work we will be mainly focusing on $b\in\left\{3,4\right\}$, which corresponds to Q3.5 and Q4.4, respectively.
 Gradient checkpointing
@@
 Instead, those matrices are multiplied with the activations $\mathbf{x}$ as $\bm{A}\left(\bm{B}\mathbf{x}\right)$.
-However, we do compute a product $\bm{A}\bm{B}$ in (5), and in the naive implementation of our method, this product together with the results of some intermediate computations (e.g., after rounding and clipping) will be automatically kept in memory for the backward pass, leading to increased memory usage.
+However, we do compute a product $\bm{A}\bm{B}$ in (5), and in the naïve implementation of our method, this product together with the results of some intermediate computations (e.g., after rounding and clipping) will be automatically kept in memory for the backward pass, leading to increased memory usage.
 To prevent this, we employ gradient checkpointing [10] on (5).
@@
 4 Experiments
-We assess the effectiveness of LR-QAT by conducting experiments on LLaMA-2 7B/13B [58], LLaMA-3 8B [1] and Mistral-0.1 7B [30].
-To compare to relevant literature with weight-activation quantization, we further experiment with LLaMa-1 7B [57].
-We first explore the impact of the choice of rank $r$, a downcasting operator $\varphi\left(\cdot\right)$, and the initialization of auxiliary matrices $\bm{A}$, $\bm{B}$.
-We then compare our method in terms of accuracy to standard full-model QAT, other baselines, and the related work.
-All detailed hyperparameters of our experiments are in Appendix A.
-Quantization
-We experiment with both weight-only and weight-activation quantization.
-The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’).
-We quantize all linear layers, except the classification head.
-In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14].
-Following OmniQuant [52] we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [42].
-Evaluation
-Following the previous work [18, 62, 52, 42], we evaluate quantized models by reporting the perplexity of language generation on WikiText-2 [46], using a sequence length of $2048$.
-We also report zero-shot accuracy on a set of common sense reasoning tasks including BoolQ [12], PIQA [5], Winogrande [51], ARC [13], and HellaSwag [66].
-For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
-Datasets and training
-We train on a small subset of SlimPajama [54], which is a close open-source replica of the dataset used for pretraining LLaMA models.
-For weight-only quantization results, we train for $10^{4}$ steps using batch size $32$ and a maximum sequence length of $1024$.
-For ablation studies and weight-activation quantization results, we use shorter training of $10^{3}$ steps.
-We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set (512 sequences).
-Baselines
-We compare with round-to-nearest quantization (RTN), where we set the ranges based on minimizing the $L^{p}$-norms between quantized and unquantized weights and report the best performing configuration.
-We also use that as initialization for LR-QAT.
-For weight-only quantization, we compare with GPTQ [18], AWQ [39], OmniQuant [52], and our implementation of PEQA [31], where we use symmetric weight quantization and following the same experimental setup and best RTN initialization, for a fair comparison.
-For weight-activation quantization, we compare to RTN, SmoothQuant [62] and LLM-QAT [42].
-Following [42], we compare to them in several different settings, where the weights, activations and KV cache values are quantized to different levels (denoted as W-A-KV).
-4.1 The impact of rank $r$
 {floatrow}
 […]
-$\varphi\left(\cdot\right)$ dtype $\bm{A}$, $\bm{B}$ init. W4 @ pc W3 @ pc
+$\varphi\!\left(\cdot\right)$ dtype $\bm{A}$, $\bm{B}$ init. W4pc W3pc
 WT-2 $\downarrow$ zero-shot $\uparrow$ WT-2 $\downarrow$ zero-shot $\uparrow$
@@
 […]
+We assess the effectiveness of LR-QAT by conducting experiments on LLaMA-2 7B/13B [61], LLaMA-3 8B [1] and Mistral-0.1 7B [31].
+To compare to relevant literature with weight-activation quantization, we further experiment with LLaMa-1 7B [60].
+We first explore the impact of the choice of rank $r$, a downcasting operator $\varphi\!\left(\cdot\right)$, and the initialization of auxiliary matrices $\bm{A}$, $\bm{B}$.
+We then compare our method in terms of accuracy to standard full-model QAT, other baselines, and the related work.
+All detailed hyperparameters of our experiments are in Appendix B.
+Quantization
+We experiment with both weight-only and weight-activation quantization.
+The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’).
+We quantize all linear layers, except the classification head.
+In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14].
+Following OmniQuant [55], we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
+Evaluation
+Following the previous work [18, 65, 55, 44], we evaluate quantized models by reporting the perplexity of language generation on WikiText-2 [48], using a sequence length of $2048$.
+We also report zero-shot accuracy on a set of common sense reasoning tasks including BoolQ [12], PIQA [5], Winogrande [54], ARC [13], and HellaSwag [69].
+For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
+Datasets and training
+We apply our methods to all linear layers in the attention blocks (both in self-attention and in the feed-forward network).
+We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$.
+We train on a small subset of SlimPajama [57], which is an open-source dataset similar to the original one used for pretraining LLaMA models.
+In all experiments we train using batch size $32$ and a maximum sequence length of $1024$.
+For weight-only and weight-activation quantization results, we train for $10^{4}$ steps.
+For ablation studies, we use shorter training of $10^{3}$ steps.
+We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set (512 sequences).
+Baselines
+We compare with round-to-nearest quantization (RTN), where we set the ranges based on minimizing the $L^{p}$-norms between quantized and unquantized weights and report the best performing configuration.
+We also use that as initialization for LR-QAT.
+For weight-only quantization, we compare with GPTQ [18], AWQ [41], OmniQuant [55], and our implementation of PEQA [32], where we use symmetric weight quantization and following the same experimental setup and best RTN initialization, for a fair comparison.
+Similar to our method, we also apply PEQA only to linear layers in the attention blocks and keep token embeddings, final classification head and LayerNorm parameters frozen.
+For weight-activation quantization, we compare our method with RTN, SmoothQuant [65], LLM-QAT [44], Outlier Suppression$+$ [63], OmniQuant [55], and our implementation of PEQA [32].
+Following [44], we compare to them in several different settings, where the weights, activations and KV cache values are quantized to different bitwidths (denoted as W-A-KV).
+[…]
+Method GPU memory, GB WikiText-2 perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
+Full-model QAT 71.9 5.74 69.29
+LR-QAT 20.5 5.68 69.43
+4.1 The impact of rank $r$
 We investigate the effect of different values of rank $r$ of the auxiliary matrices $\bm{A}$ and $\bm{B}$ and present results in Figure 2.
@@
 Going forward, we use $r=32$ in all our experiments [^2] This amounts to only 1.2% of the total number of parameters for 7B LLaMA model..
-4.2 The impact of downcasting operator $\varphi\left(\cdot\right)$ and initialization
+4.2 The choice of the downcasting operator $\varphi\!\left(\cdot\right)$ and $\bm{A}$, $\bm{B}$ initialization
 We study the effect of several choices of the downcasting operators discussed in Section 3 and summarize results in Table 2.
@@
 from FP32 to BF16, where the latter is what is commonly used for LLMs.
-4.3 Comparison with full-model QAT
-Table 3: Comparison of the proposed method with the vanilla (full-model) QAT on LLaMA-2 7B with W4 per-channel quantization.
-Method GPU memory, GB WikiText-2 perplexity $\downarrow$ Zero-shot avg. accuracy $\uparrow$
-Full model QAT 71.9 5.74 69.29
-LR-QAT 20.5 5.68 69.43
-[…]
-4.4 Main results
+Going forward, we use $\varphi={\text{{\small{Q$b$.$(8-b)$}}}}$ with ‘LoRA’ initialization and BF16 compute data type.
 Table 4: Weight-only quantization results for LLaMA-2/3 and Mistral models.
-We report WikiText-2 test set perplexity (lower is better) and zero-shot average accuracy (higher is better).
-Models marked ‘L2’, ‘L3’, and ‘M’ denote LLaMA-2, LLaMA-3, and Mistral, respectively.
+We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better).
+Models marked ‘L2’/‘L3’, and ‘M’ denote LLaMA-2/3 and Mistral, respectively.
 Numbers marked in bold are SOTA or on par (within $0.05$).
 §Uses asymmetric weight quantization.
-Bitwitdth Method WikiText-2 Perplexity $\downarrow$ Zero-shot avg. accuracy $\uparrow$
+# Bits Method WikiText-2 Perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
 L2-7B L2-13B L3-8B M-7B L2-7B L2-13B L3-8B M-7B
 FP16 5.47 4.88 6.14 5.25 70.47 73.18 74.22 75.69
-W4 @ pc RTN 6.14 5.21 7.53 5.91 68.88 71.73 72.19 73.44
+W4pc RTN 6.14 5.21 7.53 5.91 68.88 71.73 72.19 73.44
 GPTQ§ 5.83 5.13 - - - - - -
@@
 LR-QAT (ours) 5.66 5.04 6.78 5.46 69.95 73.13 73.35 73.67
-W3 @ pc RTN 26.73 8.71 34.10 9.49 43.87 55.01 47.46 64.58
+W3pc RTN 26.73 8.71 34.10 9.49 43.87 55.01 47.46 64.58
 GPTQ§ 8.37 6.44 - - - - - -
@@
 LR-QAT (ours) 6.13 5.54 8.12 6.03 67.66 71.22 70.46 71.87
-W4 @ g128 RTN 5.78 5.04 6.96 5.49 69.75 72.94 72.04 75.07
+W4g128 RTN 5.78 5.04 6.96 5.49 69.75 72.94 72.30 75.07
 GPTQ§ 5.61 4.98 - - - - - -
@@
 LR-QAT (ours) 5.59 4.97 6.62 5.36 69.88 72.89 73.72 75.18
-W3 @ g128 RTN 7.61 6.14 15.11 6.77 63.20 66.81 57.74 69.35
+W3g128 RTN 7.61 6.20 15.11 6.77 63.20 67.60 57.74 69.35
 GPTQ§ 6.29 5.42 - - - - - -
@@
 LR-QAT (ours) 5.99 5.53 7.74 5.81 67.98 70.54 70.48 71.44
+4.3 Comparison with full-model QAT
+[…]
+4.4 Main results
 Weight-only quantization
@@
 Numbers marked in bold are SOTA.
-*Evaluated with a maximum sequence length of 1024.
-Bitwitdth (W-A-KV) Method WikiText-2 $\downarrow$ Zero-shot avg. accuracy $\uparrow$
+§Uses asymmetric weight quantization.
+*Uses a maximum sequence length of 1024 for evaluation.
+# Bits (W-A-KV) Method WikiText-2 $\downarrow$ Zero-shot accuracy $\uparrow$
 BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
 FP16 5.68 75.05 79.16 70.01 72.85 44.80 76.21 69.68
-4-8-8 RTN 3.4e4 45.87 49.13 48.46 25.51 28.50 26.01 37.25
-SmoothQuant - 71.00 76.00 66.00 67.40 42.80 67.80 65.17
+4-8-8 RTN 23.18 60.55 67.14 53.43 54.08 33.11 50.77 53.18
+SmoothQuant 13.7* 71.00 76.00 66.00 67.40 42.80 67.80 65.17
 LLM-QAT 11.2* 74.60 77.50 67.70 70.20 45.60 73.50 68.18
-LR-QAT (ours) 5.97 72.97 77.86 68.98 69.91 41.30 74.35 67.56
-4-8-4 RTN 3.5e4 40.37 49.46 49.88 26.18 28.50 26.11 36.75
-SmoothQuant - 54.70 55.40 51.50 43.90 27.70 38.90 45.35
+PEQA (our impl.) 5.89 74.86 78.24 70.01 70.12 42.83 75.14 68.53
+LR-QAT (ours) 5.85 73.76 78.51 71.19 71.09 41.81 75.10 68.58
+4-8-4 RTN 182.7 48.01 52.23 48.78 29.00 22.70 28.52 38.21
+SmoothQuant 163.6* 54.70 55.40 51.50 43.90 27.70 38.90 45.35
 LLM-QAT 11.6* 69.50 75.40 64.60 66.00 43.80 69.20 64.75
-LR-QAT (ours) 6.01 75.60 78.62 69.93 69.74 41.04 74.51 68.24
+PEQA (our impl.) 6.15 72.97 77.80 67.72 67.13 40.27 73.35 66.54
+LR-QAT (ours) 6.07 73.64 77.91 67.56 69.28 41.30 73.25 67.16
+4-4-4 RTN 4.1e4 41.41 49.24 51.30 27.19 28.84 26.30 37.38
+SmoothQuant 25.25 49.10 49.80 48.00 30.40 25.80 27.40 38.42
+LLM-QAT - 61.30 51.50 51.90 27.90 23.90 31.10 41.27
+LLM-QAT + SQ - 62.40 55.90 50.60 35.50 26.40 47.80 46.43
+OS+ - 60.21 62.73 52.96 39.98 30.29 44.39 48.43
+OmniQuant§ 11.26 63.51 66.15 53.43 45.20 31.14 56.44 52.65
+PEQA (our impl.) 8.60 65.69 72.31 59.83 56.52 34.22 61.79 58.39
+LR-QAT (ours) 8.47 67.16 71.76 59.59 58.42 34.73 62.34 59.00
 We present our results for weight-activation quantization applied to LLaMA-1 7B in Table 5.
-LR-QAT consistently outperforms the PTQ baselines and is on par or better compared to LLM-QAT.
-This demonstrates that the proposed method is readily applicable not only to weight-only quantization but also weight-activation quantization, a setting that allows for a very efficient inference using fixed-point arithmetic.
+LR-QAT demonstrates superior performance compared to all the PTQ and QAT baselines, consistently outperforming them across all the bitwidth settings.
+In addition to that, as we decrease the activation bitwidths, the improvement in model performance compared to prior work becomes more pronounced.
+This indicates LR-QAT’s versatility, being readily applicable not only to weight-only quantization but also weight-activation quantization, a setting that allows for a very efficient inference using fixed-point arithmetic.
 Finally, our method can still be combined with most of the related PTQ methods including OmniQuant that shift the difficulty of activation quantization to weight quantization, and will likely lead to even better results.
@@
 Impact
-As LR-QAT improves LLM quantization and makes them therefore more efficient, we expect mainly positive outcomes of our work.
+As LR-QAT improves memory and inference efficiency of LLMs, we expect mainly positive outcomes from our work.
 Efficiently deploying LLMs will help with reducing their high power consumption at inference time.
@@
 We would like to thank Tijmen Blankevoort, Paul Whatmough, Jorn Peters, and Ties van Rozendaal for valuable discussions and support.
-Appendix A Experimental details
-In this section, we list the details related to hyperparameters and other settings used in our experiments. If not stated otherwise, the standard hyperparameters that we use are the following:
-Hyperparameter Value
+Appendix A Extended results
+In this section, we provide extended results and present some additional ablation studies.
+Table A1: LM-eval weight-only quantization results for LLaMA-2/3 and Mistral models. We report zero-shot accuracy of 6 tasks (higher is better).
+Model # Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
+LLaMA-2 7B FP16 77.74 79.11 69.14 74.58 46.25 75.98 70.47
+W4pc RTN 76.36 78.07 68.19 71.21 44.80 74.65 68.88
+PEQA (our impl.) 77.49 78.24 69.61 70.96 43.52 75.54 69.23
+LR-QAT (ours) 77.43 78.45 69.69 73.15 45.48 75.51 69.95
+W3pc RTN 46.27 60.28 54.85 38.05 23.29 40.47 43.87
+PEQA (our impl.) 71.62 76.82 66.14 65.66 39.76 72.63 65.44
+LR-QAT (ours) 74.43 77.15 68.03 69.95 43.09 73.29 67.66
+W4g128 RTN 76.76 78.18 69.77 72.60 45.73 75.43 69.75
+PEQA (our impl.) 76.88 78.89 69.85 72.18 44.11 75.95 69.64
+LR-QAT (ours) 76.73 78.62 70.48 72.85 44.97 75.62 69.88
+W3g128 RTN 66.42 75.57 65.19 64.90 38.14 68.96 63.20
+PEQA (our impl.) 75.38 77.97 68.59 70.62 42.32 73.74 68.10
+LR-QAT (ours) 73.30 78.07 67.72 71.46 43.77 73.53 67.98
+LLaMA-2 13B FP16 80.55 80.52 72.22 77.44 48.98 79.38 73.18
+W4pc RTN 79.30 79.71 70.01 75.51 48.89 76.96 71.73
+PEQA (our impl.) 78.99 80.14 71.27 76.43 48.98 79.24 72.51
+LR-QAT (ours) 79.91 80.25 71.03 77.48 50.68 79.45 73.13
+W3pc RTN 55.05 71.06 54.22 56.19 32.25 61.27 55.01
+PEQA (our impl.) 74.28 78.67 69.06 74.87 45.99 76.00 69.81
+LR-QAT (ours) 78.62 79.49 72.61 73.99 45.56 77.05 71.22
+W4g128 RTN 81.10 79.82 72.38 76.73 49.06 78.52 72.94
+PEQA (our impl.) 80.28 80.63 71.74 76.14 48.38 79.62 72.80
+LR-QAT (ours) 80.73 79.92 71.98 76.60 48.81 79.31 72.89
+W3g128 RTN 74.65 76.93 69.14 70.16 42.66 72.06 67.60
+PEQA (our impl.) 78.56 78.73 69.85 73.61 44.28 76.69 70.29
+LR-QAT (ours) 79.69 78.78 68.98 73.99 44.88 76.89 70.54
+LLaMA-3 8B FP16 81.44 80.79 72.85 77.74 53.33 79.16 74.22
+W4pc RTN 79.02 78.56 72.85 75.97 49.32 77.44 72.19
+PEQA (our impl.) 79.57 78.67 72.93 77.19 51.11 77.25 72.79
+LR-QAT (ours) 80.09 80.41 73.88 76.89 50.51 78.30 73.35
+W3pc RTN 58.65 61.75 56.04 39.60 23.81 44.91 47.46
+PEQA (our impl.) 62.63 63.66 56.59 42.34 27.13 49.35 50.28
+LR-QAT (ours) 77.46 78.51 69.85 74.83 47.35 74.73 70.46
+W4g128 RTN 79.48 79.27 73.56 75.08 48.81 77.61 72.30
+PEQA (our impl.) 80.98 80.14 72.61 76.18 49.57 78.45 72.99
+LR-QAT (ours) 80.28 80.96 74.11 77.19 51.45 78.31 73.72
+W3g128 RTN 65.47 68.39 65.19 54.00 33.45 59.96 57.74
+PEQA (our impl.) 72.26 76.06 67.80 69.02 46.08 71.89 67.19
+LR-QAT (ours) 72.97 79.38 71.67 74.37 49.06 75.44 70.48
+Mistral 7B FP16 83.58 82.10 73.88 79.59 53.92 81.07 75.69
+W4pc RTN 81.22 80.63 72.53 76.77 50.09 79.41 73.44
+PEQA (our impl.) 81.80 81.12 72.61 77.23 50.17 79.43 73.73
+LR-QAT (ours) 80.12 81.83 73.40 76.73 49.66 80.25 73.67
+W3pc RTN 68.13 77.64 63.93 63.93 41.13 72.73 64.58
+PEQA (our impl.) 80.03 80.09 69.93 72.90 45.82 77.32 71.02
+LR-QAT (ours) 81.62 80.09 70.96 74.75 46.08 77.71 71.87
+W4g128 RTN 84.16 81.77 74.43 77.95 51.71 80.42 75.07
+PEQA (our impl.) 80.89 81.72 73.80 75.42 48.46 79.76 73.34
+LR-QAT (ours) 83.85 81.77 73.88 78.54 52.22 80.79 75.18
+W3g128 RTN 78.44 79.60 69.14 71.17 43.00 74.75 69.35
+PEQA (our impl.) 81.99 81.18 69.61 74.92 47.18 78.37 72.21
+LR-QAT (ours) 81.38 80.41 70.32 73.15 45.39 78.00 71.44
+Table A2: A comparison between min-max and the best range setting used for round-to-nearest (RTN) initialization for LLaMA-2/3 and Mistral models.
+We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better).
+Model # Bits RTN init. WikiText-2 $\downarrow$ Zero-shot accuracy $\uparrow$
+LLaMA-2 7B FP16 5.47 70.47
+W4pc min-max 7.14 66.41
+best ($L^{3.5}$) 6.14 68.88
+W3pc min-max 1.9e4 35.71
+best ($L^{3.5}$) 26.73 43.87
+W4g128 best = min-max 5.78 69.75
+W3g128 min-max 8.22 64.07
+best ($L^{4}$) 7.61 63.20
+LLaMA-2 13B FP16 4.88 73.18
+W4pc min-max 5.40 72.19
+best ($L^{3.5}$) 5.21 71.73
+W3pc min-max 2.3e3 37.85
+best ($L^{5}$) 8.71 55.01
+W4g128 best = min-max 5.04 72.94
+W3g128 min-max 6.14 66.81
+best ($L^{5}$) 6.20 67.60
+LLaMA-3 8B FP16 6.14 74.22
+W4pc min-max 10.53 67.44
+best ($L^{3.5}$) 7.53 72.19
+W3pc min-max 1.6e5 35.78
+best ($L^{3.5}$) 34.10 47.46
+W4g128 min-max 6.99 72.95
+best ($L^{4}$) 6.96 72.30
+W3g128 min-max 29.38 54.54
+best ($L^{5}$) 15.11 57.74
+Mistral 7B FP16 5.25 75.69
+W4pc min-max 6.33 71.84
+best ($L^{4}$) 5.91 73.44
+W3pc min-max 3.2e3 36.78
+best ($L^{4}$) 9.49 64.58
+W4g128 min-max 5.51 74.98
+best ($L^{5}$) 5.49 75.07
+W3g128 min-max 7.22 68.35
+best ($L^{5}$) 6.77 69.35
+Appendix B Experimental details
+In this section, we list the details related to hyperparameters and other settings used in our experiments.
+If not stated otherwise, the standard hyperparameters that we use are the one shown in Table B1.
+Hyperparameter Value / Search space
 Optimizer AdamW
-LR for $\bm{A}$, $\bm{B}$ $10^{-2}$
-LR Schedule for $\bm{A}$, $\bm{B}$ Linear with Warmup
-LR for Quantization scales $10^{-5}$
-LR Schedule for Quantization scales Linear with Warmup
-Warmups Steps $10\%$ of Training Steps
-Weight Decay $0.0$
-Weight Decay for $\bm{A}$, $\bm{B}$ $0.0$
-Adam $\beta_{1}$ $0.90$
+Learning rate for $\bm{A}$, $\bm{B}$$\{10^{-5},10^{-4},10^{-3},10^{-2}\}$
+Learning rate for quantization scale ($\mathbf{s}$) $\{0{{\textsuperscript{*}}},10^{-5}\}$
+Learning rate schedule for $\bm{A}$, $\bm{B}$ linear (with warmup)
+Learning rate schedule for quantization scale ($\mathbf{s}$) linear (with warmup)
+Weight decay for $\bm{A}$, $\bm{B}$ 00
+Weight decay for quantization scale ($\mathbf{s}$) 00
+Adam $\beta_{1}$ $0.9$
 Adam $\beta_{2}$ $0.95$
 Training steps $10^{4}$
-Gradient Clipping (Maximum $L^{2}$-norm) $1.0$
-Table 6: Common hyperparameters used for experiments.
+Warmup steps $10\%$ of Training steps
+Batch size $32$
+Maximum sequence length (during training) $1024$
+$L^{2}$-norm gradient clipping (maximum norm) $1.0$
+$\alpha$ in (8) $1.0$
+Table B1: Common hyperparameters used for experiments. *Is equivalent to freezing the quantization scale obtained after initial range estimation ($\mathbf{s}=\mathbf{s_{0}}$).
 Quantization
@@
 Following OmniQuant [55] we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
+Libraries
+We implement our method in PyTorch [53] and use training and evaluation pipelines from HuggingFace libraries [20, 38, 64]. For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
 Datasets and training
-For all our experiments we use the linear with warmup learning rate schedule from HugginFace’s transformers package [61].
-To optimize the learnable parameters, we use AdamW optimizer [43] with weight decay set to zero, $\beta_{1}=0.9$ and $\beta_{2}=0.95$.
-We use a separate learning rate for quantization scales and for low-rank adapters.
-Quantization scales learning rate is always set to $10^{-5}$, where the learning rate for low-rank adapters is tuned depending on the experiment.
-For all the weight-only quantization experiments we set to $10^{-2}$, except W4g128 where we found that a smaller learning rate of $10^{-4}$ or $10^{-5}$ is beneficial.
-For weight-activation quantization experiments we set the learning rate to $10^{-5}$.
+To optimize the learnable parameters, we use AdamW optimizer [45] with weight decay set to zero, $(\beta_{1},\beta_{2})=(0.9,0.95)$ and linear learning rate warm up over the first $10\%$ steps, following by a linear decay to zero by the end of training.
+We use a separate maximum learning rate for quantization scales and for low-rank adapters, which are tuned depending on the experiment.
+We apply our methods to all linear layers in the attention blocks (both in self-attention and in the feed-forward network).
+We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$.
+Specifically, we freeze the token embeddings, the final classification heads and LayerNorm parameters.
 We train on a small subset of SlimPajama [57], which is a close open-source replica of the dataset used for pre-training LLaMA models.
-In all experiments we set $\alpha$ in (8) to $1$.
 We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set [^4] Specifically, we use the English subset of Wiki40b, https://huggingface.co/datasets/wiki40b, that contains cleaned-up text of English Wikipedia and training/validation splits.
 (512 sequences).
-For weight-only quantization results we train for $10^{3}$ steps, but for LLaMA-v3 and Mistral we only train for $5000$ steps due to constrained resources.
-In all cases we used a number of warmup steps equal to the $10\%$ of the training steps, a batch size of $32$, and a maximum sequence length of $1024$.
-For ablation studies we always use shorter training of $1000$ steps, and for weight-activation quantization results we use training of $1000$ steps.
-Baselines
+For weight-only and weight-activation quantization results, we train for $10^{4}$ steps (but for LLaMA-3 and Mistral we only train for $5000$ steps due to constrained resources).
+For ablation studies, we always use shorter training of $10^{3}$ steps.
+The rest of the hyperparameters and their search spaces are listed in Table B1
+PTQ initialization
 We compare with vanilla round-to-nearest quantization (RTN), where we explore several choices of range setting and report the best one based on Wikipedia validation set perplexity, and also use that as initialization for our method.

### v3

@@
 Unlike most related work, our method (i) is inference-efficient, leading to no additional overhead compared to traditional PTQ; (ii) can be seen as a general extended pretraining framework, meaning that the resulting model can still be utilized for any downstream task afterwards; (iii) can be applied across a wide range of quantization settings, such as different choices quantization granularity, activation quantization, and seamlessly combined with many PTQ techniques.
-We apply LR-QAT to LLaMA-2/3 and Mistral model families and validate its effectiveness on several downstream tasks.
+We apply LR-QAT to LLaMA-1/2/3 and Mistral model families and validate its effectiveness on several downstream tasks.
 Our method outperforms common post-training quantization (PTQ) approaches and reaches the same model performance as full-model QAT at the fraction of its memory usage.
 Specifically, we can train a 7B LLM on a single consumer grade GPU with 24GB of memory.
+Our source code is available at https://github.com/qualcomm-ai-research/LR-QAT.
 1 Introduction
@@
 Inspired by parameter-efficient fine-tuning (PEFT) and low-rank adaptation (LoRA) literature, we propose Low-Rank Quantization-Aware Training (LR-QAT) – a lightweight memory-efficient and inference-efficient QAT algorithm for LLMs.
-LR-QAT reduces the memory requirements of training a 7B LLM from >70GB of GPU memory to <21GB without degrading the predictive performance compared to traditional full-model QAT, making it possible to train such models on a single consumer grade GPU.
+LR-QAT reduces the memory requirements of training a 7B LLM from >98GB of GPU memory to <21GB without degrading the predictive performance compared to traditional full-model QAT, making it possible to train such models on a single consumer grade GPU.
 Unlike most related work that combines low-rank adaptation with quantization, our method is also inference-efficient.
@@
 (3) Finally, we combine the proposed quantization formulation with gradient checkpointing to avoid aggressive memory spikes from storing some of the intermediate results in memory for the backward pass.
-We apply LR-QAT to LLaMA-2/3 and Mistral model families and demonstrate its effectiveness on several general language modeling datasets and zero-shot evaluation on some of the common reasoning downstream tasks.
+We apply LR-QAT to LLaMA-1/2/3 and Mistral model families and demonstrate its effectiveness on several general language modeling datasets and zero-shot evaluation on some of the common reasoning downstream tasks.
 Our method outperforms recent PTQ approaches and reaches the same predictive performance as full-model QAT at the fraction of its memory usage.
@@
 $\mathbf{x}$ and $\mathbf{y}$ denote the input and the output of the linear layer.
-Right: Memory requirements for training with various QAT techniques on LLaMA-2 7B (assuming batch size 1, sequence length 1024, $r=32$, and BF16 compute data type).
+Right: Memory requirements for training with various QAT techniques on LLaMA-2 7B, assuming batch size 1, sequence length 1024, $r=32$, and BF16 compute data type.
+‘Intermediate results’ refer to the results of some intermediate computations, e.g.
+after rounding/clipping in (3), which are saved in memory for the backward pass.
 Neural network quantization is one of the most powerful ways to reduce model footprint, data transfer and compute requirements.
@@
 QAT with low-rank adapters
-Let’s recall how traditional QAT works.
+Let’s recall how traditional QAT [17] works.
 Given a linear layer with a weight matrix $\bm{W}\in\mathbb{R}^{m\times k}$ and assuming $b$-bit symmetric uniform affine quantization, the quantization is simulated as follows:
@@
 Gradient checkpointing
-Note that both in the original LoRA paper [25] and in the related work like QLoRA [16], there is no need to compute the product $\bm{A}\bm{B}$.
+Note that both in the original LoRA paper [25] and in the related work like QLoRA [16], there is no need to compute the product $\bm{A}\bm{B}$ explicitly.
 Instead, those matrices are multiplied with the activations $\mathbf{x}$ as $\bm{A}\left(\bm{B}\mathbf{x}\right)$.
-However, we do compute a product $\bm{A}\bm{B}$ in (5), and in the naïve implementation of our method, this product together with the results of some intermediate computations (e.g., after rounding and clipping) will be automatically kept in memory for the backward pass, leading to increased memory usage.
+However, we have to compute a product $\bm{A}\bm{B}$ in (5), and in the naïve implementation of our method, this product together with the results of some intermediate computations (e.g., after rounding and clipping) will be automatically kept in memory for the backward pass, leading to increased memory usage.
 To prevent this, we employ gradient checkpointing [10] on (5).
@@
 4 Experiments
-{floatrow}
-[…]
-$\varphi\!\left(\cdot\right)$ dtype $\bm{A}$, $\bm{B}$ init. W4pc W3pc
-WT-2 $\downarrow$ zero-shot $\uparrow$ WT-2 $\downarrow$ zero-shot $\uparrow$
-FP32 FP32 LoRA 5.69 69.28 6.21 66.62
-FP16 FP32 LoRA $+$0.00 --0.13 $+$0.01 --0.01
-BF16 FP32 LoRA --0.01 $+$0.11 $+$0.01 $+$0.45
-Q4.4 / Q3.5 FP32 LoRA --0.01 $+$0.16 $+$0.01 $+$0.31
-Q4.4 / Q3.5 BF16 LoRA --0.01 $+$0.15 $+$0.01 $+$0.31
-INT-4 / INT-3 FP32 LoRA $+$0.02 --0.04 $+$1.5e3 --30.30
-INT-4 / INT-3 FP32 LoftQ ($T=1$) $+$0.14 --1.02 $+$0.18 $+$0.26
-INT-4 / INT-3 FP32 LoftQ ($T=64$) $+$0.40 --1.40 $+$1.37 --2.01
-[…]
-[figure]
-[…]
-We assess the effectiveness of LR-QAT by conducting experiments on LLaMA-2 7B/13B [61], LLaMA-3 8B [1] and Mistral-0.1 7B [31].
-To compare to relevant literature with weight-activation quantization, we further experiment with LLaMa-1 7B [60].
+We assess the effectiveness of LR-QAT by conducting experiments on LLaMA 7B [60], LLaMA-2 7B/13B [61], LLaMA-3 8B [1], and Mistral-0.1 7B [31].
 We first explore the impact of the choice of rank $r$, a downcasting operator $\varphi\!\left(\cdot\right)$, and the initialization of auxiliary matrices $\bm{A}$, $\bm{B}$.
@@
 We experiment with both weight-only and weight-activation quantization.
-The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’).
+The default settings are INT4 / INT3 / INT2 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’).
+We use symmetric quantization, except the INT2 case, where we use asymmetric quantization (1), for a fair comparison with related work.
 We quantize all linear layers, except the classification head.
@@
 Following OmniQuant [55], we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
+Datasets and training
+We apply our method to all linear layers in the attention blocks, both in self-attention and in the feed-forward network.
+We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$ and keep embedding layers, final classification head and RMSNorm parameters frozen.
+In the case of asymmetric weight quantization, a zero offset $\mathbf{z}$ is set during range estimation phase and kept frozen throughout training ($\mathbf{z}=\mathbf{z_{0}}$).
+We train on a small subset of SlimPajama [57], which is an open-source dataset similar to the original one used for pretraining LLaMA models.
+In all experiments we train using batch size $32$ and a maximum sequence length of $1024$.
+For all weight-only and weight-activation quantization results, we train for $10^{4}$ steps.
+For ablation studies in Sections 4.1 and 4.2 we use shorter training of $10^{3}$ steps.
+We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set (512 sequences).
 Evaluation
@@
 For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
-Datasets and training
-We apply our methods to all linear layers in the attention blocks (both in self-attention and in the feed-forward network).
-We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$.
-We train on a small subset of SlimPajama [57], which is an open-source dataset similar to the original one used for pretraining LLaMA models.
-In all experiments we train using batch size $32$ and a maximum sequence length of $1024$.
-For weight-only and weight-activation quantization results, we train for $10^{4}$ steps.
-For ablation studies, we use shorter training of $10^{3}$ steps.
-We select hyperparameters based on the perplexity of a small subset of Wikipedia validation set (512 sequences).
+Specifically, we use lm_eval v0.4.2 and report acc_norm for tasks where it’s available (PIQA, ARC-e, ARC-c, HellaSwag) and otherwise acc (BoolQ and Winogrande).
 Baselines
@@
 We also use that as initialization for LR-QAT.
-For weight-only quantization, we compare with GPTQ [18], AWQ [41], OmniQuant [55], and our implementation of PEQA [32], where we use symmetric weight quantization and following the same experimental setup and best RTN initialization, for a fair comparison.
-Similar to our method, we also apply PEQA only to linear layers in the attention blocks and keep token embeddings, final classification head and LayerNorm parameters frozen.
+For weight-only quantization, we compare with GPTQ [18], AWQ [41], and OmniQuant [55].
+We also compare with our implementation of PEQA [32] and full-model QAT (LSQ) [17], where we follow the same experimental setup as for our method, together with the best RTN initialization, for a fair comparison.
 For weight-activation quantization, we compare our method with RTN, SmoothQuant [65], LLM-QAT [44], Outlier Suppression$+$ [63], OmniQuant [55], and our implementation of PEQA [32].
 Following [44], we compare to them in several different settings, where the weights, activations and KV cache values are quantized to different bitwidths (denoted as W-A-KV).
+4.1 The impact of rank $r$
+{floatrow}
 […]
-Method GPU memory, GB WikiText-2 perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
-Full-model QAT 71.9 5.74 69.29
-LR-QAT 20.5 5.68 69.43
-4.1 The impact of rank $r$
+$\varphi\!\left(\cdot\right)$ dtype $\bm{A}$, $\bm{B}$ init. WikiText-2 $\downarrow$ Zero-shot acc. $\uparrow$
+W4 pc W3 pc W4 pc W3 pc
+FP32 FP32 LoRA 5.69 6.21 69.28 66.62
+FP16 FP32 LoRA $+$0.00 $+$0.01 --0.13 --0.01
+BF16 FP32 LoRA --0.01 $+$0.01 $+$0.11 $+$0.45
+Q4.4 / Q3.5 FP32 LoRA --0.01 $+$0.01 $+$0.16 $+$0.31
+Q4.4 / Q3.5 BF16 LoRA --0.01 $+$0.01 $+$0.15 $+$0.31
+INT-4 / INT-3 FP32 LoRA $+$0.02 $+$20.5 --0.04 --22.8
+INT-4 / INT-3 FP32 LoftQ ($T=1$) $+$0.28 $+$0.18 --0.67 $+$0.26
+INT-4 / INT-3 FP32 LoftQ ($T=64$) $+$0.40 $+$1.37 --1.40 --2.01
+[…]
+[figure]
+[…]
 We investigate the effect of different values of rank $r$ of the auxiliary matrices $\bm{A}$ and $\bm{B}$ and present results in Figure 2.
@@
 Even though $\operatorname{rank}\!\left(\bm{A}\bm{B}\right)=1$, by applying a low-rank approximation inside the rounding and clipping operators in (8), this can overall lead to a high-rank perturbation to the original weights $\bm{\Phi_{0}}$ (in the integer domain).
+Finally, for all ranks we observe only a small standard deviation between $0.005$ and $0.008$ ppl., indicating the robustness of LR-QAT to a random initialization of $\bm{B}$.
 Going forward, we use $r=32$ in all our experiments [^2] This amounts to only 1.2% of the total number of parameters for 7B LLaMA model..
@@
 Going forward, we use $\varphi={\text{{\small{Q$b$.$(8-b)$}}}}$ with ‘LoRA’ initialization and BF16 compute data type.
-Table 4: Weight-only quantization results for LLaMA-2/3 and Mistral models.
+4.3 Comparison with full-model QAT
+[…]
+Method GPU mem., Time/100 steps, WikiText-2 ppl. $\downarrow$ Zero-shot acc. $\uparrow$
+GB sec W4 pc W3 pc W4 pc W3 pc
+Full-model QAT (LSQ) 62.2 (98.5) 3248±${}^{\scriptsize\pm{}\text{7}}$ 5.77±${}^{\scriptsize\pm{}\text{0.02}}$ 6.14±${}^{\scriptsize\pm{}\text{0.01}}$ 68.96±${}^{\scriptsize\pm{}\text{0.29}}$ 67.14±${}^{\scriptsize\pm{}\text{0.13}}$
+LR-QAT (ours) 20.5 1522±${}^{\scriptsize\pm{}\text{5}}$ 5.66±${}^{\scriptsize\pm{}\text{0.00}}$ 6.13±${}^{\scriptsize\pm{}\text{0.02}}$ 69.72±${}^{\scriptsize\pm{}\text{0.32}}$ 67.70±${}^{\scriptsize\pm{}\text{0.25}}$
+[…]
+As we can see in Table 3, training with our method leads to on par or better predictive performance at a significantly lower memory usage and runtime compared to full-model QAT.
+We include results for other models and bitwidths in Table 4.
+A more detailed runtime comparison can also be found in Table A8.
+4.4 Main results
+Table 4: Weight-only quantization results for LLaMA and Mistral models.
 We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better).
-Models marked ‘L2’/‘L3’, and ‘M’ denote LLaMA-2/3 and Mistral, respectively.
+Models marked ‘L1’/‘L2’/‘L3’, and ‘M’ denote LLaMA-1/2/3 and Mistral, respectively.
 Numbers marked in bold are SOTA or on par (within $0.05$).
 §Uses asymmetric weight quantization.
+✕ denotes out of memory.
 # Bits Method WikiText-2 perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
-L2-7B L2-13B L3-8B M-7B L2-7B L2-13B L3-8B M-7B
-FP16 5.47 4.88 6.14 5.25 70.47 73.18 74.22 75.69
-W4pc RTN 6.14 5.21 7.53 5.91 68.88 71.73 72.19 73.44
-GPTQ§ 5.83 5.13 - - - - - -
-AWQ 6.15 5.12 - - - - - -
-OmniQuant§ 5.74 5.02 - - - - - -
-PEQA (our impl.) 5.71 5.03 7.51 5.56 69.23 72.51 72.79 73.73
-LR-QAT (ours) 5.66 5.04 6.78 5.46 69.95 73.13 73.35 73.67
-W3pc RTN 26.73 8.71 34.10 9.49 43.87 55.01 47.46 64.58
-GPTQ§ 8.37 6.44 - - - - - -
-AWQ 24.00 10.45 - - - - - -
-OmniQuant§ 6.58 5.58 - - - - - -
-PEQA (our impl.) 6.45 5.73 27.45 6.51 65.44 69.81 50.28 71.02
-LR-QAT (ours) 6.13 5.54 8.12 6.03 67.66 71.22 70.46 71.87
-W4g128 RTN 5.78 5.04 6.96 5.49 69.75 72.94 72.30 75.07
-GPTQ§ 5.61 4.98 - - - - - -
-AWQ 5.62 4.97 - - - - - -
-OmniQuant§ 5.58 4.95 - - - - - -
-PEQA (our impl.) 5.67 5.02 6.89 5.48 69.64 72.80 72.99 73.34
-LR-QAT (ours) 5.59 4.97 6.62 5.36 69.88 72.89 73.72 75.18
-W3g128 RTN 7.61 6.20 15.11 6.77 63.20 67.60 57.74 69.35
-GPTQ§ 6.29 5.42 - - - - - -
-AWQ 6.24 5.32 - - - - - -
-OmniQuant§ 6.03 5.28 - - - - - -
-PEQA (our impl.) 6.05 5.58 9.64 5.85 68.10 70.29 67.19 72.21
-LR-QAT (ours) 5.99 5.53 7.74 5.81 67.98 70.54 70.48 71.44
-4.3 Comparison with full-model QAT
-[…]
-4.4 Main results
+L1-7B L2-7B L2-13B L3-8B M-7B L1-7B L2-7B L2-13B L3-8B M-7B
+FP16 5.68 5.47 4.88 6.14 5.25 69.68 70.47 73.18 74.22 75.69
+W4 pc RTN 6.33 6.14 5.21 7.53 5.91 68.51 68.88 71.73 72.19 73.44
+GPTQ§ 6.13 5.83 5.13 - - 64.95 - - - -
+AWQ 6.08 6.15 5.12 - - - - - - -
+OmniQuant§ 5.86 5.74 5.02 - - - - - - -
+LSQ (our impl.) 5.94 5.77 6.87 5.73 68.37 68.96 73.28 72.88
+PEQA (our impl.) 5.86 5.71 5.03 7.51 5.56 68.49 69.23 72.51 72.79 73.73
+LR-QAT (ours) 5.84 5.66 5.03 6.78 5.46 68.54 69.72 73.19 73.84 74.44
+W4 g128 RTN 6.05 5.78 5.04 6.96 5.49 68.93 69.75 72.94 72.30 75.07
+GPTQ§ 5.85 5.61 4.98 - - - - - - -
+AWQ 5.81 5.62 4.97 - - - - - - -
+OmniQuant§ 5.77 5.58 4.95 - - - - - - -
+LSQ (our impl.) 5.76 5.61 6.58 5.67 69.17 69.68 73.31 72.90
+PEQA (our impl.) 5.75 5.67 5.02 6.89 5.48 69.19 69.64 72.80 72.99 73.34
+LR-QAT (ours) 5.75 5.59 4.97 6.57 5.37 69.15 69.88 72.91 73.66 75.28
+W3 pc RTN 12.88 26.73 8.71 34.10 9.49 54.66 43.87 55.01 47.46 64.58
+GPTQ§ 8.06 8.37 6.44 - - - - - - -
+AWQ 11.88 24.00 10.45 - - - - - - -
+OmniQuant§ 6.49 6.58 5.58 - - - - - - -
+LSQ (our impl.) 6.29 6.14 8.14 6.06 66.29 67.14 69.58 71.61
+PEQA (our impl.) 6.56 6.45 5.73 26.20 6.51 65.75 65.44 69.81 51.05 71.02
+LR-QAT (ours) 6.27 6.13 5.54 8.12 6.03 66.60 67.70 71.22 70.46 71.87
+W3 g128 RTN 7.96 7.61 6.20 15.11 6.77 63.50 63.20 67.60 57.74 69.35
+GPTQ§ 6.55 6.29 5.42 - - - - - - -
+AWQ 6.46 6.24 5.32 - - - - - - -
+OmniQuant§ 6.15 6.03 5.28 - - - - - - -
+LSQ (our impl.) 6.20 6.02 8.08 5.90 66.53 68.36 70.11 71.96
+PEQA (our impl.) 6.22 6.05 5.58 9.64 5.85 66.66 68.10 70.29 67.19 72.21
+LR-QAT (ours) 6.17 5.98 5.32 7.74 5.80 66.81 68.62 71.51 70.48 72.41
+W2 pc RTN§ 4.9e3 5.2e3 5.2e3 6.4e4 6.8e3 37.92 36.52 36.27 36.80 36.59
+GPTQ§ 2.1e3 7.7e3 2.1e3 - - - - - - -
+OmniQuant§ 15.47 37.37 17.21 - - - - - - -
+PEQA (our impl.)§ 8.24 9.34 7.51 14.47 8.08 59.83 57.40 62.29 57.72 64.48
+LR-QAT (ours)§ 7.99 8.51 7.16 12.52 7.98 61.77 60.03 65.28 58.49 65.11
+W2 g128 RTN§ 708 2.5e3 115.6 1.4e4 369 39.74 37.94 41.12 36.97 41.30
+GPTQ§ 44.01 36.77 28.14 - - - - - - -
+AWQ 2.6e5 2.2e5 1.2e5 - - - - - - -
+OmniQuant§ 9.72 11.06 8.26 - - - - - - -
+PEQA (our impl.)§ 8.12 7.87 6.83 12.39 8.05 60.69 60.74 64.74 57.81 64.65
+LR-QAT (ours)§ 7.86 7.62 6.57 11.09 7.92 61.60 61.70 66.75 60.46 65.27
 Weight-only quantization
@@
 As we can see, in almost most cases LR-QAT outperforms or is on par with prior weight-only quantization methods across various LLM families and quantization settings, including both per-channel and group-wise quantization.
+In the case of extremely low-bitwidth regime, our method consistently and significantly outperforms related work, across all settings.
 In a few cases, especially in case of group-wise quantization our method did not outpeform OmniQuant.
@@
 Weight-activation quantization
-Table 5: Weight and activation quantization results for LLaMA-1 7B.
-We report WikiText-2 test set perplexity (lower is better) and zero-shot accuracy of 6 tasks (higher is better).
+Table 5: Weight and activation quantization results for LLaMA-1/2 (denoted ‘L1’/‘L2’, respectively).
+We report WikiText-2 test set perplexity and zero-shot accuracy of 6 tasks.
 Numbers marked in bold are SOTA.
@@
 *Uses a maximum sequence length of 1024 for evaluation.
-# Bits (W-A-KV) Method WikiText-2 $\downarrow$ Zero-shot accuracy $\uparrow$
-BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
-FP16 5.68 75.05 79.16 70.01 72.85 44.80 76.21 69.68
-4-8-8 RTN 23.18 60.55 67.14 53.43 54.08 33.11 50.77 53.18
-SmoothQuant 13.7* 71.00 76.00 66.00 67.40 42.80 67.80 65.17
-LLM-QAT 11.2* 74.60 77.50 67.70 70.20 45.60 73.50 68.18
-PEQA (our impl.) 5.89 74.86 78.24 70.01 70.12 42.83 75.14 68.53
-LR-QAT (ours) 5.85 73.76 78.51 71.19 71.09 41.81 75.10 68.58
-4-8-4 RTN 182.7 48.01 52.23 48.78 29.00 22.70 28.52 38.21
-SmoothQuant 163.6* 54.70 55.40 51.50 43.90 27.70 38.90 45.35
-LLM-QAT 11.6* 69.50 75.40 64.60 66.00 43.80 69.20 64.75
-PEQA (our impl.) 6.15 72.97 77.80 67.72 67.13 40.27 73.35 66.54
-LR-QAT (ours) 6.07 73.64 77.91 67.56 69.28 41.30 73.25 67.16
-4-4-4 RTN 4.1e4 41.41 49.24 51.30 27.19 28.84 26.30 37.38
-SmoothQuant 25.25 49.10 49.80 48.00 30.40 25.80 27.40 38.42
-LLM-QAT - 61.30 51.50 51.90 27.90 23.90 31.10 41.27
-LLM-QAT + SQ - 62.40 55.90 50.60 35.50 26.40 47.80 46.43
-OS+ - 60.21 62.73 52.96 39.98 30.29 44.39 48.43
-OmniQuant§ 11.26 63.51 66.15 53.43 45.20 31.14 56.44 52.65
-PEQA (our impl.) 8.60 65.69 72.31 59.83 56.52 34.22 61.79 58.39
-LR-QAT (ours) 8.47 67.16 71.76 59.59 58.42 34.73 62.34 59.00
-We present our results for weight-activation quantization applied to LLaMA-1 7B in Table 5.
-LR-QAT demonstrates superior performance compared to all the PTQ and QAT baselines, consistently outperforming them across all the bitwidth settings.
+# Bits (W-A-KV) Method WikiText-2 perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
+L1-7B L2-7B L2-13B L1-7B L2-7B L2-13B
+FP16 5.68 5.47 4.88 69.68 70.47 73.18
+4-8-8 RTN 6.88 6.17 5.23 65.83 68.55 71.54
+SmoothQuant 13.7* - - 65.17 - -
+LLM-QAT 11.2* - - 68.18 - -
+PEQA (our impl.) 5.89 5.72 5.08 68.53 69.11 72.49
+LR-QAT (ours) 5.85 5.67 5.04 68.58 69.32 73.18
+4-8-4 RTN 7.66 6.85 5.78 62.78 65.16 67.06
+SmoothQuant 163.6* - - 45.35 - -
+LLM-QAT 11.6* - - 64.75 - -
+PEQA (our impl.) 6.15 6.03 5.27 66.54 67.56 71.35
+LR-QAT (ours) 6.07 5.90 5.24 67.16 69.84 71.65
+4-4-4 RTN 17.75 18.98 11.37 49.60 51.75 55.07
+SmoothQuant 25.25 83.12 35.88 38.42 - -
+LLM-QAT - - - 41.27 - -
+LLM-QAT + SQ - - - 46.43 - -
+Outlier Suppression+ - - - 48.43 - -
+OmniQuant§ 11.26 14.26 12.30 52.65 - -
+PEQA (our impl.) 8.60 8.72 7.23 58.39 57.93 62.26
+LR-QAT (ours) 8.47 8.46 7.15 59.00 58.98 62.65
+We present our results for weight-activation quantization applied to LLaMA-1/2 models in Table 5.
+LR-QAT demonstrates superior performance compared to all the PTQ and QAT baselines, consistently outperforming them across all model families and the bitwidth settings.
 In addition to that, as we decrease the activation bitwidths, the improvement in model performance compared to prior work becomes more pronounced.
 This indicates LR-QAT’s versatility, being readily applicable not only to weight-only quantization but also weight-activation quantization, a setting that allows for a very efficient inference using fixed-point arithmetic.
-Finally, our method can still be combined with most of the related PTQ methods including OmniQuant that shift the difficulty of activation quantization to weight quantization, and will likely lead to even better results.
+Further, our method can still be combined with most of the related PTQ methods including OmniQuant that shift the difficulty of activation quantization to weight quantization, and will likely lead to even better results.
 5 Discussion
@@
 Appendix A Extended results
-In this section, we provide extended results and present some additional ablation studies.
-Table A1: LM-eval weight-only quantization results for LLaMA-2/3 and Mistral models. We report zero-shot accuracy of 6 tasks (higher is better).
-Model # Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
-LLaMA-2 7B FP16 77.74 79.11 69.14 74.58 46.25 75.98 70.47
+In this section, we provide additional detailed results.
+Table A1: LM-eval weight-only quantization results for LLaMA-1 7B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
+# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
+FP16 75.05 79.16 70.01 72.85 44.80 76.21 69.68
+W4 pc RTN 73.18 78.78 69.14 71.38 44.37 74.22 68.51
+GPTQ§ 67.70 76.00 66.70 66.90 43.00 69.40 64.95
+LSQ (our impl.) 73.91 78.24 69.22 70.84 43.26 74.74 68.37
+PEQA (our impl.) 74.71 78.29 70.09 70.33 42.24 75.27 68.49
+LR-QAT (ours) 74.13 78.29 70.01 71.21 42.41 75.16 68.54
+W4 g128 RTN 74.77 78.51 70.64 71.30 43.60 74.74 68.93
+LSQ (our impl.) 75.90 79.22 70.01 71.42 43.34 75.15 69.17
+PEQA (our impl.) 75.75 79.17 70.17 70.75 43.60 75.71 69.19
+LR-QAT (ours) 75.29 78.62 69.61 71.59 44.11 75.67 69.15
+W3 pc RTN 58.93 70.40 55.72 55.01 32.17 55.75 54.66
+LSQ (our impl.) 71.35 77.97 68.82 66.33 40.10 73.14 66.29
+PEQA (our impl.) 72.69 77.15 65.90 68.27 38.91 71.60 65.75
+LR-QAT (ours) 73.24 78.18 67.40 67.47 40.53 72.77 66.60
+W3 g128 RTN 69.48 76.33 64.40 64.44 38.65 67.67 63.50
+LSQ (our impl.) 71.04 77.97 68.11 68.27 40.44 73.37 66.53
+PEQA (our impl.) 71.65 78.24 68.51 68.18 40.10 73.30 66.66
+LR-QAT (ours) 72.84 78.02 67.40 68.52 41.04 73.04 66.81
+W2 pc RTN§ 43.24 52.61 51.62 27.48 26.28 26.26 37.92
+PEQA (our impl.)§ 67.19 74.32 61.80 59.97 32.42 63.28 59.83
+LR-QAT (ours)§ 68.07 74.27 64.80 61.03 36.95 65.48 61.77
+W2 g128 RTN§ 41.16 58.05 50.20 34.97 23.12 30.93 39.74
+PEQA (our impl.)§ 67.40 74.21 62.04 60.98 35.32 64.16 60.69
+LR-QAT (ours)§ 67.65 74.86 61.80 62.37 37.03 65.90 61.60
+Table A2: LM-eval weight-only quantization results for LLaMA-2 7B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
+# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
+FP16 77.74 79.11 69.14 74.58 46.25 75.98 70.47
 W4 pc RTN 76.36 78.07 68.19 71.21 44.80 74.65 68.88
+LSQ (our impl.) 76.69 78.15 68.12 71.84 44.18 74.76 68.96
 PEQA (our impl.) 77.49 78.24 69.61 70.96 43.52 75.54 69.23
-LR-QAT (ours) 77.43 78.45 69.69 73.15 45.48 75.51 69.95
-W3pc RTN 46.27 60.28 54.85 38.05 23.29 40.47 43.87
-PEQA (our impl.) 71.62 76.82 66.14 65.66 39.76 72.63 65.44
-LR-QAT (ours) 74.43 77.15 68.03 69.95 43.09 73.29 67.66
+LR-QAT (ours) 77.41 78.56 69.42 72.80 44.66 75.45 69.72
 W4 g128 RTN 76.76 78.18 69.77 72.60 45.73 75.43 69.75
+LSQ (our impl.) 77.28 78.45 69.61 72.18 44.97 75.61 69.68
 PEQA (our impl.) 76.88 78.89 69.85 72.18 44.11 75.95 69.64
 LR-QAT (ours) 76.73 78.62 70.48 72.85 44.97 75.62 69.88
+W3 pc RTN 46.27 60.28 54.85 38.05 23.29 40.47 43.87
+LSQ (our impl.) 74.39 77.91 66.85 69.15 41.53 73.04 67.14
+PEQA (our impl.) 71.62 76.82 66.14 65.66 39.76 72.63 65.44
+LR-QAT (ours) 75.08 77.73 67.50 69.97 42.70 73.24 67.70
 W3 g128 RTN 66.42 75.57 65.19 64.90 38.14 68.96 63.20
+LSQ (our impl.) 74.68 77.80 68.35 71.76 44.11 73.46 68.36
 PEQA (our impl.) 75.38 77.97 68.59 70.62 42.32 73.74 68.10
-LR-QAT (ours) 73.30 78.07 67.72 71.46 43.77 73.53 67.98
-LLaMA-2 13B FP16 80.55 80.52 72.22 77.44 48.98 79.38 73.18
+LR-QAT (ours) 76.61 77.31 68.98 72.05 42.58 74.20 68.62
+W2 pc RTN§ 39.57 50.71 50.20 26.73 26.28 25.63 36.52
+PEQA (our impl.)§ 64.98 71.33 58.09 55.22 33.11 61.66 57.40
+LR-QAT (ours)§ 69.88 73.07 62.90 56.57 34.04 63.71 60.03
+W2 g128 RTN§ 39.08 55.93 50.99 29.84 24.06 27.71 37.94
+PEQA (our impl.)§ 69.51 73.83 63.46 58.21 34.90 64.53 60.74
+LR-QAT (ours)§ 70.18 73.83 63.69 60.27 34.98 67.24 61.70
+Table A3: LM-eval weight-only quantization results for LLaMA-2 13B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization. ✕ denotes out of memory.
+# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
+FP16 80.55 80.52 72.22 77.44 48.98 79.38 73.18
 W4 pc RTN 79.30 79.71 70.01 75.51 48.89 76.96 71.73
+LSQ (our impl.)
 PEQA (our impl.) 78.99 80.14 71.27 76.43 48.98 79.24 72.51
-LR-QAT (ours) 79.91 80.25 71.03 77.48 50.68 79.45 73.13
+LR-QAT (ours) 80.15 80.09 72.06 77.65 49.91 79.28 73.19
+W4 g128 RTN 81.10 79.82 72.38 76.73 49.06 78.52 72.94
+LSQ (our impl.)
+PEQA (our impl.) 80.28 80.63 71.74 76.14 48.38 79.62 72.80
+LR-QAT (ours) 80.73 80.30 71.74 76.14 49.06 79.51 72.91
 W3 pc RTN 55.05 71.06 54.22 56.19 32.25 61.27 55.01
+LSQ (our impl.)
 PEQA (our impl.) 74.28 78.67 69.06 74.87 45.99 76.00 69.81
 LR-QAT (ours) 78.62 79.49 72.61 73.99 45.56 77.05 71.22
-W4g128 RTN 81.10 79.82 72.38 76.73 49.06 78.52 72.94
-PEQA (our impl.) 80.28 80.63 71.74 76.14 48.38 79.62 72.80
-LR-QAT (ours) 80.73 79.92 71.98 76.60 48.81 79.31 72.89
 W3 g128 RTN 74.65 76.93 69.14 70.16 42.66 72.06 67.60
+LSQ (our impl.)
 PEQA (our impl.) 78.56 78.73 69.85 73.61 44.28 76.69 70.29
-LR-QAT (ours) 79.69 78.78 68.98 73.99 44.88 76.89 70.54
-LLaMA-3 8B FP16 81.44 80.79 72.85 77.74 53.33 79.16 74.22
+LR-QAT (ours) 79.79 79.60 70.64 74.24 46.76 78.00 71.51
+W2 pc RTN§ 38.35 48.97 48.54 27.78 27.99 25.97 36.27
+PEQA (our impl.)§ 71.19 74.48 59.67 62.96 37.29 68.14 62.29
+LR-QAT (ours)§ 72.60 76.61 66.22 66.67 39.76 69.79 65.28
+W2 g128 RTN§ 50.12 57.29 50.36 34.43 22.18 32.35 41.12
+PEQA (our impl.)§ 70.34 76.88 66.85 64.94 38.74 70.66 64.74
+LR-QAT (ours)§ 75.47 77.86 65.98 67.63 41.04 72.50 66.75
+Table A4: LM-eval weight-only quantization results for LLaMA-3 8B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
+# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
+FP16 81.44 80.79 72.85 77.74 53.33 79.16 74.22
 W4 pc RTN 79.02 78.56 72.85 75.97 49.32 77.44 72.19
+LSQ (our impl.) 80.58 80.03 72.69 78.24 50.17 77.94 73.28
 PEQA (our impl.) 79.57 78.67 72.93 77.19 51.11 77.25 72.79
-LR-QAT (ours) 80.09 80.41 73.88 76.89 50.51 78.30 73.35
+LR-QAT (ours) 81.62 79.98 72.85 78.32 52.05 78.19 73.84
+W4 g128 RTN 79.48 79.27 73.56 75.08 48.81 77.61 72.30
+LSQ (our impl.) 80.24 80.36 73.01 77.23 50.43 78.60 73.31
+PEQA (our impl.) 80.98 80.14 72.61 76.18 49.57 78.45 72.99
+LR-QAT (ours) 80.40 80.90 73.48 77.44 51.11 78.60 73.66
 W3 pc RTN 58.65 61.75 56.04 39.60 23.81 44.91 47.46
-PEQA (our impl.) 62.63 63.66 56.59 42.34 27.13 49.35 50.28
+LSQ (our impl.) 76.42 78.24 70.01 72.31 45.22 75.26 69.58
+PEQA (our impl.) 63.18 64.74 57.62 43.39 26.88 50.48 51.05
 LR-QAT (ours) 77.46 78.51 69.85 74.83 47.35 74.73 70.46
-W4g128 RTN 79.48 79.27 73.56 75.08 48.81 77.61 72.30
-PEQA (our impl.) 80.98 80.14 72.61 76.18 49.57 78.45 72.99
-LR-QAT (ours) 80.28 80.96 74.11 77.19 51.45 78.31 73.72
 W3 g128 RTN 65.47 68.39 65.19 54.00 33.45 59.96 57.74
+LSQ (our impl.) 75.50 78.78 69.69 73.57 48.55 74.55 70.11
 PEQA (our impl.) 72.26 76.06 67.80 69.02 46.08 71.89 67.19
 LR-QAT (ours) 72.97 79.38 71.67 74.37 49.06 75.44 70.48
-Mistral 7B FP16 83.58 82.10 73.88 79.59 53.92 81.07 75.69
+W2 pc RTN§ 44.89 48.97 47.51 25.29 27.73 26.41 36.80
+PEQA (our impl.)§ 62.54 72.96 59.51 56.82 33.70 60.77 57.72
+LR-QAT (ours)§ 65.81 72.31 64.72 53.54 33.45 61.11 58.49
+W2 g128 RTN§ 38.47 53.32 51.78 28.75 22.70 26.81 36.97
+PEQA (our impl.)§ 58.53 72.52 61.96 57.41 35.32 61.14 57.81
+LR-QAT (ours)§ 67.98 73.83 62.83 56.94 36.43 64.77 60.46
+Table A5: LM-eval weight-only quantization results for Mistral 7B. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
+# Bits Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
+FP16 83.58 82.10 73.88 79.59 53.92 81.07 75.69
 W4 pc RTN 81.22 80.63 72.53 76.77 50.09 79.41 73.44
+LSQ (our impl.) 81.96 80.41 73.09 75.08 48.55 78.19 72.88
 PEQA (our impl.) 81.80 81.12 72.61 77.23 50.17 79.43 73.73
-LR-QAT (ours) 80.12 81.83 73.40 76.73 49.66 80.25 73.67
+LR-QAT (ours) 81.99 81.28 73.56 78.20 51.02 80.57 74.44
+W4 g128 RTN 84.16 81.77 74.43 77.95 51.71 80.42 75.07
+LSQ (our impl.) 80.40 80.63 73.16 76.26 48.72 78.24 72.90
+PEQA (our impl.) 80.89 81.72 73.80 75.42 48.46 79.76 73.34
+LR-QAT (ours) 83.55 81.61 74.51 78.28 52.90 80.84 75.28
 W3 pc RTN 68.13 77.64 63.93 63.93 41.13 72.73 64.58
+LSQ (our impl.) 79.97 80.74 70.24 74.37 47.18 77.14 71.61
 PEQA (our impl.) 80.03 80.09 69.93 72.90 45.82 77.32 71.02
 LR-QAT (ours) 81.62 80.09 70.96 74.75 46.08 77.71 71.87
-W4g128 RTN 84.16 81.77 74.43 77.95 51.71 80.42 75.07
-PEQA (our impl.) 80.89 81.72 73.80 75.42 48.46 79.76 73.34
-LR-QAT (ours) 83.85 81.77 73.88 78.54 52.22 80.79 75.18
 W3 g128 RTN 78.44 79.60 69.14 71.17 43.00 74.75 69.35
+LSQ (our impl.) 81.01 80.36 71.03 74.41 47.27 77.68 71.96
 PEQA (our impl.) 81.99 81.18 69.61 74.92 47.18 78.37 72.21
-LR-QAT (ours) 81.38 80.41 70.32 73.15 45.39 78.00 71.44
-Table A2: A comparison between min-max and the best range setting used for round-to-nearest (RTN) initialization for LLaMA-2/3 and Mistral models.
+LR-QAT (ours) 81.71 80.90 70.48 75.08 47.78 78.50 72.41
+W2 pc RTN§ 39.42 51.03 49.49 26.22 26.71 26.67 36.59
+PEQA (our impl.)§ 74.34 75.95 66.22 62.21 38.48 69.68 64.48
+LR-QAT (ours)§ 75.08 76.39 65.98 64.48 39.59 69.14 65.11
+W2 g128 RTN§ 54.98 56.91 51.70 31.27 23.21 29.70 41.30
+PEQA (our impl.)§ 75.41 75.84 64.96 65.45 38.23 68.00 64.65
+LR-QAT (ours)§ 75.81 77.15 64.64 63.59 39.51 70.89 65.27
+Table A6: LM-eval weight and activation quantization results for LLaMA models. We report zero-shot accuracy of 6 tasks (higher is better). §Uses asymmetric weight quantization.
+Model # Bits (W-A-KV) Method BoolQ PIQA Winogrande ARC-e ARC-c HellaSwag Avg.
+LLaMA-1 7B FP16 75.05 79.16 70.01 72.85 44.80 76.21 69.68
+4-8-8 RTN 71.35 76.66 66.46 66.84 41.55 72.10 65.83
+SmoothQuant 71.00 76.00 66.00 67.40 42.80 67.80 65.17
+LLM-QAT 74.60 77.50 67.70 70.20 45.60 73.50 68.18
+PEQA (our impl.) 74.86 78.24 70.01 70.12 42.83 75.14 68.53
+LR-QAT (ours) 73.76 78.51 71.19 71.09 41.81 75.10 68.58
+4-8-4 RTN 68.81 75.46 62.12 62.46 39.51 68.33 62.78
+SmoothQuant 54.70 55.40 51.50 43.90 27.70 38.90 45.35
+LLM-QAT 69.50 75.40 64.60 66.00 43.80 69.20 64.75
+PEQA (our impl.) 72.97 77.80 67.72 67.13 40.27 73.35 66.54
+LR-QAT (ours) 73.64 77.91 67.56 69.28 41.30 73.25 67.16
+4-4-4 RTN 50.49 64.25 52.41 48.27 30.12 52.04 49.60
+SmoothQuant 49.10 49.80 48.00 30.40 25.80 27.40 38.42
+LLM-QAT 61.30 51.50 51.90 27.90 23.90 31.10 41.27
+LLM-QAT + SQ 62.40 55.90 50.60 35.50 26.40 47.80 46.43
+Outlier Suppression+ 60.21 62.73 52.96 39.98 30.29 44.39 48.43
+OmniQuant§ 63.51 66.15 53.43 45.20 31.14 56.44 52.65
+PEQA (our impl.) 65.69 72.31 59.83 56.52 34.22 61.79 58.39
+LR-QAT (ours) 67.16 71.76 59.59 58.42 34.73 62.34 59.00
+LLaMA-2 7B FP16 77.74 79.11 69.14 74.58 46.25 75.98 70.47
+4-8-8 RTN 75.87 77.91 67.88 71.09 44.03 74.51 68.55
+PEQA (our impl.) 77.37 77.97 69.77 70.54 43.52 75.50 69.11
+LR-QAT (ours) 77.00 78.13 69.14 72.10 44.11 75.44 69.32
+4-8-4 RTN 70.37 76.01 63.38 68.94 41.47 70.76 65.16
+PEQA (our impl.) 74.71 77.48 67.40 69.28 42.75 73.75 67.56
+LR-QAT (ours) 74.46 77.69 68.51 69.78 42.75 73.82 67.84
+4-4-4 RTN 57.86 64.91 54.46 49.62 31.83 51.83 51.75
+PEQA (our impl.) 67.09 70.67 60.06 54.80 32.17 62.76 57.93
+LR-QAT (ours) 66.94 71.98 60.77 57.20 33.87 63.10 58.98
+LLaMA-2 13B FP16 80.55 80.52 72.22 77.44 48.98 79.38 73.18
+4-8-8 RTN 79.24 79.27 70.01 75.51 48.29 76.92 71.54
+PEQA (our impl.) 79.02 80.20 71.19 76.60 48.72 79.22 72.49
+LR-QAT (ours) 81.10 79.76 71.35 77.36 50.60 78.91 73.18
+4-8-4 RTN 77.25 76.61 66.69 67.72 41.13 72.98 67.06
+PEQA (our impl.) 78.01 79.22 69.30 75.59 48.21 77.78 71.35
+LR-QAT (ours) 78.59 79.54 70.80 75.29 48.04 77.64 71.65
+4-4-4 RTN 62.60 67.90 53.20 57.32 34.13 55.25 55.07
+PEQA (our impl.) 68.72 74.21 62.35 63.85 36.01 68.44 62.26
+LR-QAT (ours) 70.64 73.88 63.14 61.78 38.14 68.31 62.65
+Table A7: A comparison between min-max and the best range setting used for round-to-nearest (RTN) initialization for LLaMA and Mistral models.
 We report WikiText-2 test set perplexity (lower is better) and average zero-shot accuracy (higher is better).
-Model # Bits RTN init. WikiText-2 $\downarrow$ Zero-shot accuracy $\uparrow$
-LLaMA-2 7B FP16 5.47 70.47
-W4pc min-max 7.14 66.41
-best ($L^{3.5}$) 6.14 68.88
-W3pc min-max 1.9e4 35.71
-best ($L^{3.5}$) 26.73 43.87
-W4g128 best = min-max 5.78 69.75
-W3g128 min-max 8.22 64.07
-best ($L^{4}$) 7.61 63.20
-LLaMA-2 13B FP16 4.88 73.18
-W4pc min-max 5.40 72.19
-best ($L^{3.5}$) 5.21 71.73
-W3pc min-max 2.3e3 37.85
-best ($L^{5}$) 8.71 55.01
-W4g128 best = min-max 5.04 72.94
-W3g128 min-max 6.14 66.81
-best ($L^{5}$) 6.20 67.60
-LLaMA-3 8B FP16 6.14 74.22
-W4pc min-max 10.53 67.44
-best ($L^{3.5}$) 7.53 72.19
-W3pc min-max 1.6e5 35.78
-best ($L^{3.5}$) 34.10 47.46
-W4g128 min-max 6.99 72.95
-best ($L^{4}$) 6.96 72.30
-W3g128 min-max 29.38 54.54
-best ($L^{5}$) 15.11 57.74
-Mistral 7B FP16 5.25 75.69
-W4pc min-max 6.33 71.84
-best ($L^{4}$) 5.91 73.44
-W3pc min-max 3.2e3 36.78
-best ($L^{4}$) 9.49 64.58
-W4g128 min-max 5.51 74.98
-best ($L^{5}$) 5.49 75.07
-W3g128 min-max 7.22 68.35
-best ($L^{5}$) 6.77 69.35
+§Uses asymmetric weight quantization.
+# Bits Range WikiText-2 perplexity $\downarrow$ Avg. zero-shot accuracy $\uparrow$
+estimator L1-7B L2-7B L2-13B L3-8B M-7B L1-7B L2-7B L2-13B L3-8B M-7B
+FP16 5.68 5.47 4.88 6.14 5.25 69.68 70.47 73.18 74.22 75.69
+W4 pc best est. $L^{4}$ $L^{3.5}$ $L^{3.5}$ $L^{3.5}$ $L^{4}$ $L^{4}$ $L^{3.5}$ $L^{3.5}$ $L^{3.5}$ $L^{4}$
+best 6.33 6.14 5.21 7.53 5.91 68.51 68.88 71.73 72.19 73.44
+min-max 6.85 7.14 5.40 10.53 6.33 66.23 66.41 72.19 67.44 71.84
+W4 g128 best est. $L^{5}$ min-max min-max $L^{4}$ $L^{5}$ $L^{5}$ min-max min-max $L^{4}$ $L^{5}$
+best 6.05 5.78 5.04 6.96 5.49 68.93 69.75 72.94 72.30 75.07
+min-max 6.08 5.78 5.04 6.99 5.51 68.96 69.75 72.94 72.95 74.98
+W3 pc best est. $L^{3.5}$ $L^{3.5}$ $L^{5}$ $L^{3.5}$ $L^{4}$ $L^{3.5}$ $L^{3.5}$ $L^{5}$ $L^{3.5}$ $L^{4}$
+best 12.88 26.73 8.71 34.10 9.49 54.66 43.87 55.01 47.46 64.58
+min-max 2.4e4 1.9e4 2.3e3 1.6e5 3.2e3 36.02 35.71 37.85 35.78 36.78
+W3 g128 best est. $L^{5}$ $L^{4}$ $L^{5}$ $L^{5}$ $L^{5}$ $L^{5}$ $L^{4}$ $L^{5}$ $L^{5}$ $L^{5}$
+best 7.95 7.61 6.20 15.11 6.77 63.50 63.20 67.60 57.74 69.35
+min-max 8.10 8.22 6.14 29.38 7.22 62.69 64.07 66.81 54.54 68.35
+W2 pc§ best est. $L^{2.4}$ $L^{3}$ $L^{3}$ $L^{3.5}$ $L^{3.5}$ $L^{2.4}$ $L^{3}$ $L^{3}$ $L^{3.5}$ $L^{3.5}$
+best 4.9e3 5.2e3 5.2e3 6.4e4 6.8e3 37.92 36.52 36.27 36.80 36.59
+min-max 1.1e5 2.5e4 4.9e4 1.4e6 7.5e4 39.07 36.53 39.01 37.69 38.13
+W2 g128§ best est. $L^{4}$ $L^{3.5}$ $L^{5}$ $L^{4}$ $L^{5}$ $L^{4}$ $L^{3.5}$ $L^{5}$ $L^{4}$ $L^{5}$
+best 708 2.5e3 115.6 1.4e4 369 39.74 37.94 41.12 36.97 41.30
+min-max 3.6e3 5.9e3 341 2.8e5 3.4e3 37.06 36.78 40.30 37.18 37.16
+[…]
+[…]
+grad. accumulation steps
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Appendix B Experimental details
@@
 Optimizer AdamW
-Learning rate for $\bm{A}$, $\bm{B}$$\{10^{-5},10^{-4},10^{-3},10^{-2}\}$
-Learning rate for quantization scale ($\mathbf{s}$) $\{0{{\textsuperscript{*}}},10^{-5}\}$
+Learning rate for $\bm{A}$, $\bm{B}$ (INT4 / INT3 / INT2) $\{10^{-5},10^{-4},10^{-3},10^{-2}\}$
+Learning rate for $\mathbf{s}$ (INT4 / INT3) $\{0{{\textsuperscript{*}}},10^{-5}\}$
+Learning rate for $\mathbf{s}$ (INT2) $\{10^{-5},10^{-4}\}$
+[…]
+Learning rate for $\bm{W}$ (full-model QAT only) $\{1,5,10,50,100\}\cdot 10^{-5}$
 Learning rate schedule for $\bm{A}$, $\bm{B}$ linear (with warmup)
-Learning rate schedule for quantization scale ($\mathbf{s}$) linear (with warmup)
+Learning rate schedule for $\mathbf{s}$ linear (with warmup)
+Learning rate schedule for $\bm{W}$ (full-model QAT only) linear (with warmup)
 Weight decay for $\bm{A}$, $\bm{B}$ 00
-Weight decay for quantization scale ($\mathbf{s}$) 00
-Adam $\beta_{1}$ $0.9$
-Adam $\beta_{2}$ $0.95$
+Weight decay for $\mathbf{s}$ 00
+Weight decay for $\bm{W}$ (full-model QAT only) $0.1$
+Adam $\left(\beta_{1},\beta_{2}\right)$ $\left(0.9,0.95\right)$
 Training steps $10^{4}$
@@
 $\alpha$ in (8) $1.0$
-Table B1: Common hyperparameters used for experiments. *Is equivalent to freezing the quantization scale obtained after initial range estimation ($\mathbf{s}=\mathbf{s_{0}}$).
+Table B1: Common hyperparameters used for experiments.
+*Is equivalent to freezing the quantization scale / zero offset obtained after initial range estimation ($\mathbf{s}=\mathbf{s_{0}}$, $\mathbf{z}=\mathbf{z_{0}}$).
 Quantization
 We experiment with both weight-only and weight-activation quantization.
-The default settings are INT4 / INT3 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’).
-We always use symmetric uniform affine quantization.
+The default settings are INT4 / INT3 / INT2 per-channel (denoted ‘pc’) and group-wise weight quantization with a group size of 128 (denoted ‘g128’).
+We use symmetric quantization, except the INT2 case, where we use asymmetric quantization (1), for a fair comparison with related work.
 We quantize all linear layers, except the classification head.
-LayerNorm and Embedding layers are always kept at full precision.
 In weight-activation quantization, defaults are INT4 per-channel weight and per-token activation quantization [14].
-Following OmniQuant [55] we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
+Following OmniQuant [55], we quantize all inputs to matmuls with exception of the softmax output and additionally quantize the KV-cache as in LLM-QAT [44].
 Libraries
-We implement our method in PyTorch [53] and use training and evaluation pipelines from HuggingFace libraries [20, 38, 64]. For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
+We implement our method in PyTorch [53] and use training and evaluation pipelines from HuggingFace libraries [20, 38, 64].
+For zero-shot evaluation, we use the LM Evaluation Harness framework [19].
+Specifically, we use lm_eval v0.4.2 and report acc_norm for tasks where it’s available (PIQA, ARC-e, ARC-c, HellaSwag) and otherwise acc (BoolQ and Winogrande).
 Datasets and training
@@
 We only train low-rank auxiliary matrices $\bm{A}$, $\bm{B}$ and the quantization parameters $\mathbf{s}$.
-Specifically, we freeze the token embeddings, the final classification heads and LayerNorm parameters.
+Specifically, we freeze embedding layers, the final classification heads and RMSNorm parameters.
+In the case of asymmetric weight quantization, a zero offset $\mathbf{z}$ is set during range estimation phase and kept frozen throughout training.
 We train on a small subset of SlimPajama [57], which is a close open-source replica of the dataset used for pre-training LLaMA models.
@@
 (512 sequences).
-For weight-only and weight-activation quantization results, we train for $10^{4}$ steps (but for LLaMA-3 and Mistral we only train for $5000$ steps due to constrained resources).
-For ablation studies, we always use shorter training of $10^{3}$ steps.
-The rest of the hyperparameters and their search spaces are listed in Table B1
+For all weight-only and weight-activation quantization results, including the comparison with full-model QAT in Section 4.3, we train for $10^{4}$ steps.
+For ablation studies in Sections 4.1 and 4.2 we use shorter training of $10^{3}$ steps.
+Since the full-model QAT experiment requires more than 80GB of GPU memory, we apply checkpointing on the quantization function $\widehat{\bm{W}}$ to be able to run the experiment on an Nvidia A100 80GB GPU.
+We have also experimented with CPU optimizer state offloading, but that turned out to be significantly slower, see details in Table A8.
+The rest of the hyperparameters and their search spaces are listed in Table B1.
 PTQ initialization
@@
 Computational Resources
-Each experiment for which we reported results, was executed on a single NVidia A100 GPU equipped with 80GB of VRAM.
-LLaMA-2 7B and 13B experiments needed respectively around 3 and 5 days for 10k training steps experiments.
-Mistral 7B and LLaMA-3 8B needed around 1.6 days for 5k training steps experiments.
-For obtaining all the results in the paper, including the ablations, we needed 107 GPU days (A100).
-Including preliminary experiments that did not make it in the final paper and hyperparameter turning we estimate the total compute costs of this research to approximately 500 GPU days.
+All the experiments were executed on a single Nvidia A100 GPU equipped with 80GB of VRAM.
+Models of size 7B, 8B and 13B needed respectively around 2.25, 2.75 and 4.5 days for $10^{4}$ training steps experiments.
+For obtaining all the results in the paper, including the ablations, we needed 202 GPU days (A100).
+Including preliminary experiments that did not make it in the final paper and hyperparameter turning we estimate the total compute costs of this research to approximately 750 GPU days.